### v3

# General Hamiltonian description of nonreciprocal interactions

Abstract
In a vast class of systems, which includes members as diverse as active colloids and bird flocks, interactions do not stem from a potential, and are in general nonreciprocal. Thus, it is not possible to define a conventional energy function, nor to use analytical or numerical tools that rely on it. Here, we overcome these limitations by constructing a Hamiltonian that includes auxiliary degrees of freedom; when subject to a constraint, this Hamiltonian yields the original non-reciprocal dynamics. We show that Glauber dynamics based on the constrained Hamiltonian reproduces the steady states of the original Langevin dynamics, as we explicitly illustrate for dissipative XY spins with vision-cone interactions. Further, the symplectic structure inherent to our construction allows us to apply the well-developed notions of Hamiltonian engineering, which we demonstrate by varying the amplitude of a periodic drive to tune the spin interactions between those of a square and a chain lattice geometry. Overall, our framework for generic nonreciprocal pairwise interactions paves the way for bringing to bear the full conceptual and methodological power of conventional statistical mechanics and Hamiltonian dynamics to nonreciprocal systems.
I Introduction
Newton’s third law states that every action has a reaction equal in magnitude and opposite in direction. For two particles interacting via a potential, the interaction forces are reciprocal: $\mathbf{F}_{2\to 1}=-\mathbf{F}_{1\to 2}$. However, for many mesoscopic and macroscopic particles, interactions do not arise from a potential but are instead mediated by a non-equilibrium field [1]. For example, interactions between colloids can be mediated by hydrodynamic flows, by a dissolved chemical, or by a temperature field [2, 3, 4, 5, 6, 7]. Respectively, interactions between animals, such as birds, wildebeest or humans, can be mediated by their vision field [8, 9, 10]. In such systems, interactions are in general nonreciprocal: $\mathbf{F}_{2\to 1}\neq-\mathbf{F}_{1\to 2}$. Nonreciprocal interactions have been reported in systems including spinning starfish embryos [11] as well as active colloids and droplets [12, 13, 14, 15, 16, 17, 18, 19, 20]; they have also been programmed in colloids [21, 22] and in robots [23, 24, 25]. Despite recent advances [26, 27, 25, 28, 29, 30, 31], a general statistical-mechanical framework for nonreciprocal systems is still missing [26].
[…]
Figure 1: Classical systems can be classified as either Hamiltonian (e.g., harmonic oscillator) or non-Hamiltonian. The latter obey forces which can either be derived from a potential (e.g., sedimenting particle) or not (e.g., nonreciprocal systems). Here we construct a Hamiltonian embedding for nonreciprocal systems.
Non-Hamiltonian systems, on the other hand, lack conjugate variables, and hence do not enjoy any of the advantages offered by the symplectic structure. Broadly speaking, they fall into two categories: systems whose equations of motion are governed by conservative forces that can be derived from a potential, and those that cannot, cf. Fig. 1. Perhaps the most common example of a non-Hamiltonian potential system is a particle subject to a conservative force, such as gravity, and coupled to a dissipative medium, such as a liquid  [^1] We define dissipation by lack of conservation of phase-space volume under the dynamics.. A notable advantage of having a potential is the ability to use Monte-Carlo methods [38] to investigate the properties of the system coupled to a thermal bath.
Conversely, in the simplest setting, nonreciprocal systems are modeled by forces that cannot be derived from a single mutual interaction potential [26]. These nonconservative forces do not admit an energy function and are, furthermore, non-Hamiltonian in general. Physically, this means that nonreciprocal systems are intrinsically out of thermal equilibrium, which significantly complicates their study beyond the bare simulation of the underlying equations of motion. For instance, due to the absence of a potential, it is a priori unclear how to apply Monte-Carlo methods to simulate and analyze properties of their non-equilibrium steady states. Similarly, due to the lack of conjugate variables, deriving effective descriptions for nonreciprocal models cannot be done using canonical transformations.
In this work, we remove these obstacles by designing a Hamiltonian embedding for nonreciprocal systems. Specifically, for any equation of motion with pairwise interactions described by nonreciprocal velocity-independent forces, we present a construction of a reciprocal Hamiltonian, comprised of the original system and an auxiliary degree of freedom for each original one [Fig. 2]. By imposing a constraint on each (original, auxiliary)-pair, the equation of motion derived from this Hamiltonian reproduces the original nonreciprocal dynamics. Our results apply equally to inertial and dissipative nonreciprocal dynamics; for dissipative nonreciprocal systems in particular, the original and auxiliary degrees of freedom form canonically conjugate variables.
We exemplify our construction using dissipative XY spins on a square lattice with vision-cone interactions — a paradigmatic example of nonreciprocal interactions [39, 29, 40, 41]. To demonstrate the utility of our construction, we first show (both analytically and numerically) the equivalence of the steady states obtained from Monte-Carlo simulations with Glauber dynamics based on the Hamiltonian embedding on the one hand, and the original equations of motion subject to Langevin noise on the other. We then use the emergent symplectic structure of the Hamiltonian embedding to analyze the vision-cone XY system subject to a periodic drive using Floquet theory [42], and derive effective equations of motion for the driven system. In the SM [43], we provide a general proof for the construction, accompanied by applications to four experimental systems featuring nonreciprocal interactions: metal-dielectric Janus particles [15, 20], walker robots [44], robotic metamaterials [23], and feedback-controlled active particles [21].
Figure 2: (a) Schematic illustration of XY spins with vision-cone interactions on a square lattice: each spin (blue arrow) described by the angle $\theta_{j}$ interacts only with nearest neighbors within its vision cone of angle $\Psi$ (cyan); a pair of unidirectional thin/thick green arrows show the nonreciprocal interaction strength. The vector from site $i$ to site $j$ is denoted by $\mathbf{r}_{ij}{=}a(\cos\psi_{ij},\sin\psi_{ij})$. (b) Corresponding Hamiltonian embedding: A layer of auxiliary spins $\varphi_{j}$ (red arrows) is coupled to the original spins (blue) reciprocally (cf. inset and Eq. (2)). Imposing the constraint that auxiliary spins mirror the original ones, $\theta_{i}(0){-}\varphi_{i}(i){=}\pi$, we recover the equations of motion for the nonreciprocal system from the equations of motion of its Hamiltonian embedding.
II XY spins with nonreciprocal Vision-Cone Interactions
Consider a square lattice of linear dimension $L$ and lattice constant $a$ with periodic boundary conditions. Each lattice site $i$ contains a classical spin whose orientation, $\mathbf{S}_{i}{=}(\cos\theta_{i},\sin\theta_{i})$, can rotate continuously within the lattice plane, and is parametrized by the angle $\theta_{i}$. The spins interact via pairwise vision-cone interactions of strength
$J_{ij}(\theta_{i})=\left\{\begin{array}[]{cc}J,&\text{for}\;\min\left\{2\pi-\left|\theta_{i}-\psi_{ij}\right|,\left|\theta_{i}-\psi_{ij}\right|\right\}\leq\Psi\\ 0,&\mathrm{else,}\end{array}\right.$
where $\Psi$ denotes half the vision cone angle, and $\psi_{ij}$ is the angle between the line connecting sites $i$ and $j$ and the horizontal [Fig. 2(a), inset]. The reciprocal XY model is recovered for $\Psi{=}\pi$. We couple the system to an external bath of temperature $T$ through a Gaussian (white) noise term $\eta_{i}(t)$ with zero mean and unit variance; this bath defines the notion of temperature used throughout this work. The interacting spins undergo Langevin dynamics, following the equation of motion
$\dot{\theta}_{i}\left(t\right)=-\sum_{j\in\{\langle ij\rangle\}}J_{ij}(\theta_{i})\sin(\theta_{i}-\theta_{j})+\sqrt{2T}\eta_{i}(t),$ (1)
where the Boltzmann constant $k_{B}{=}1$, and the damping coefficient $\gamma{=}1$. The corresponding vision-cone XY interaction [Fig. 2(a)] gives rise to a minimal model for nonreciprocal interactions, which has recently gained significant attention. Previous work has studied the role of non-reciprocity on the transition from the disordered to the ferromagnetic phase [39, 29, 40, 41], as well as on the creation and annihilation of topological defects [45, 40].
Since a spin may be in the vision cone of its neighbor, but not the other way around [see Fig. 2(a)], vision-cone interactions are nonreciprocal: $J_{ij}(\theta_{i}){\neq}J_{ji}(\theta_{j})$; hence, the magnitudes of the torques acting on spins $i$ and $j$, as given by the right-hand-side of Eq. (1), differ. Thus, one cannot assign an unambiguous interaction energy to each lattice bond [Fig. 2(a), inset], and therefore vision-cone interactions cannot be derived from a potential.
III Constrained Hamiltonian Embedding
We now present a procedure that enables us to obtain nonreciprocal interactions in the equations of motion from an effective Hamiltonian function, in the absence of the noise term $\eta_{i}$. This is achieved by first extending the model by adding an auxiliary set of degrees of freedom, i.e., doubling the configuration space. Introducing suitable couplings between original and auxiliary degrees of freedom using reciprocal interactions allows us to construct a well-defined Hamiltonian. The original nonreciprocal dynamics is generated from Hamilton’s equations of motion for states obeying one constraint on each pair of (original, auxiliary) degrees of freedom, as explained below. This constraint reduces the size of the configuration space back to that of the original nonreciprocal system. Crucially, the constraint is preserved by the dynamics, so that it suffices to impose it once, e.g. on the initial state, of the time evolution.
Let us demonstrate the procedure using the vision-cone dynamics from Eq. (1), see SI [43] for more examples. We introduce a set of auxiliary spins, $\mathbf{a}_{i}{=}(\cos\varphi_{i},\sin\varphi_{i})$, on an auxiliary lattice layer [red spins in Fig. 2(b)], which point opposite to the spins in the original system: $\mathbf{a}_{i}{=}{-}\mathbf{S}_{i}$. To grasp the essence of the idea, consider first two neighboring spins $\mathbf{S}_{i},\mathbf{S}_{j}$ which interact nonreciprocally. The key insight is to symmetrize the nonreciprocal interaction between $\mathbf{S}_{i}$ and $\mathbf{S}_{j}$, but then use the interaction with the neighboring auxiliary spins $\mathbf{a}_{i},\mathbf{a}_{j}$ to each cancel one part of the symmetric coupling between the $\mathbf{S}$-spins, cf. Fig. 2(b) inset. Note that this cancellation only works if the auxiliary spins are antiparallel to the original, $\mathbf{a}_{i}{=}{-}\mathbf{S}_{i}$. This is the origin of the constraint that we need to enforce, which we call the mirror constraint. There is no interaction between the auxiliary spins $\mathbf{a}_{i}$.
Following this procedure, we write a Hamiltonian
$H(\theta_{i},\varphi_{i})=H_{SS}(\theta_{i})+H_{Sa}(\theta_{i},\varphi_{i}),$ (2)
[…]
The first term, $H_{SS}$, describes the symmetrized interactions within the original system,
$H_{SS}=-\sum_{\langle ij\rangle}\left[J_{ij}(\theta_{i})+J_{ji}(\theta_{j})\right]\cos(\theta_{i}-\theta_{j}).$ (3)
Here, the energy of bond $\langle ij\rangle$ is given by the sum of the vision-cone interactions along that bond, $J_{\left\langle ij\right\rangle}\left(\theta_{i},\theta_{j}\right){=}J_{ij}(\theta_{i}){+}J_{ji}(\theta_{j})$. Although this symmetrized interaction is reciprocal, it depends on the orientations $\theta_{i}$ and $\theta_{j}$.
The second term, $H_{Sa}$, represents the interaction between the two layers, where the auxiliary spins $\mathbf{a}_{i}$ interact with the original spins $\mathbf{S}_{j}$, following
$H_{Sa}=\sum_{\langle ij\rangle}\left[-J_{ij}(\theta_{i})\cos(\theta_{j}-\varphi_{i})-J_{ji}(\theta_{j})\cos(\theta_{i}-\varphi_{j})\right].$
The magnitude of each of the two terms in $H_{Sa}$ above directly corresponds to that in the symmetrized interaction, see Fig. 2(b) inset. As we show next, this is essential to recover the original nonreciprocal dynamics.
The equations of motion generated by the Hamiltonian (2), $\dot{\theta}_{i}{=}\left\{\theta_{i},H\right\}$ and $\dot{\varphi}_{i}{=}\left\{\varphi_{i},H\right\}$, form a set of equations that couple the auxiliary and original degrees of freedom:
$\dot{\theta}_{i}\left(t\right)$ $=$ $\partial_{\varphi_{i}}H=\sum_{j\in\{\langle ij\rangle\}}J_{ij}\left(\theta_{i}\right)\sin\left(\varphi_{i}-\theta_{j}\right),$ (4)
$\dot{\varphi}_{i}\left(t\right)$ $=$ $-\partial_{\theta_{i}}H=\sum_{j\in\{\langle ij\rangle\}}-J_{ij}(\theta_{i})\sin(\theta_{i}-\theta_{j})$
$-J_{ji}(\theta_{j})\left[\sin(\theta_{i}-\varphi_{j})+\sin(\theta_{i}-\theta_{j})\right]$
$+\partial_{\theta_{i}}J_{ij}(\theta_{i})\left[\cos(\theta_{j}-\varphi_{i})+\cos(\theta_{j}-\theta_{i})\right].$
The $\partial_{\theta_{i}}J_{ij}(\theta_{i})$ term appears since the coupling strength $J_{ij}(\theta_{i})$ depends on the spin degree of freedom $\theta_{i}$; this implicit dependence is at the origin of the discrepancy between the dynamics generated by the reciprocal textbook XY model with the replacement $J_{ij}\mapsto J_{ij}(\theta_{i})$ and the dynamics of Eq. (1[45].
If the mirror constraint is imposed on the initial conditions,
$\mathbf{a}_{i}(0)=-\mathbf{S}_{i}(0)\quad\Longleftrightarrow\quad\theta_{i}(0)-\varphi_{i}(0)=\pi,$ (5)
it is preserved under the equations of motion, i.e., $\theta_{i}(t){-}\varphi_{i}(t){=}\pi$ for all times, as we show in the SI [43]. Because, under the constraint, the dynamics of the $\varphi_{i}(t)$ degree of freedom is tied to that of the $\theta_{i}(t)$ spins, the equations of motion (4) uncouple. Moreover, each degree of freedom recovers precisely the nonreciprocal vision-cone interaction term from Eq. (1). Formally, the constraint (5) causes the unwanted terms in the last two lines of Eq. (4) to cancel, and makes the remaining equations the same.
Note that the form of the Hamiltonian embedding is not unique. For instance, one can distribute the interaction symmetrically among the two types of degrees of freedom without changing the resulting equations of motion upon imposing the constraint. It is currently an open question whether this freedom can be used to imprint additional properties into the embedding Hamiltonian.
The procedure illustrated above is not limited to vision-cone XY interactions. In the SI [43], we first present a general proof for arbitrary velocity-independent nonreciprocal interactions. To illustrate the wide applicability of our framework, we construct the Hamiltonian for four specific examples: metal-dielectric Janus particles [20, 15], self-aligning active particles such as walker robots [44, 46], robotic metamaterials [23], and feedback-controlled active particles [21]. In these examples, the dynamics are overdamped, i.e., dissipation dominates over inertia. However, using suitable kinetic-energy terms, the Hamiltonian embedding also applies to inertial (i.e., non-dissipative) nonreciprocal dynamics, suggesting a broad applicability, see SI [43].
In the following, we illustrate the utility of the proposed Hamiltonian description by (i) demonstrating that the long-time steady state of the original Langevin dynamics (1) coincides with the steady state of Glauber dynamics generated by the Hamiltonian (2), thus enabling the use of Monte-Carlo methods to study nonreciprocal dynamics. In addition, (ii) we engineer the strength of vision cone interactions along different lattice directions with the help of high-frequency periodic drives, explicitly using the symplectic structure of the Hamiltonian formulation; this enables the application of Floquet engineering to control the phase transitions of nonreciprocal systems.
IV Equivalence between Hamiltonian Monte-Carlo and Langevin Dynamics
The absence of a well-defined interaction energy per bond precludes the direct application of Monte-Carlo methods to study nonreciprocal systems. To circumvent this problem, recent works have introduced new concepts and constructions [47, 48, 49] to perform Monte-Carlo simulations, such as the ‘selfish energy’ of a lattice site $i$, $E_{i}{=}{-}\sum_{j\in\langle ij\rangle}J_{ij}(\theta_{i})\cos(\theta_{i}{-}\theta_{j})$ [30, 29, 50], or modified acceptance rates that include nonreciprocal interactions through a Legendre-like transform similar to a grand-canonical ensemble [51]. A notable common disadvantage is that these tools are ad-hoc, i.e., it is a priori unclear whether the resulting steady states correspond to those of the physical microscopic Langevin dynamics. Additionally, they do not enjoy formal mathematical guarantees for convergence, unlike Monte-Carlo methods for reciprocal systems [38].
Starting from the reciprocal Hamiltonian in Eq. (2), we propose a two-step Monte-Carlo algorithm. First, we apply Glauber dynamics at temperature $T$ defined by the transition rates [52]
$w\left(\theta_{i}\rightarrow\theta_{i}^{\prime}\right)=\frac{1}{2}\left(1-\tanh\frac{H(\theta_{i}^{\prime},\varphi_{i})-H(\theta_{i},\varphi_{i})}{2T}\right),$ (6)
to the original $\mathbf{S}$-spin layer while keeping the auxiliary spin degree of freedom $\varphi_{j}$ fixed. In a second step, we enforce the constraint by deterministically adjusting the auxiliary spins according to the accepted move: $\varphi_{i}{\rightarrow}\varphi^{\prime}_{i}{=}\theta_{i}^{\prime}{+}\pi$; this causes the sampler to converge to a non-equilibrium steady state. Our complete algorithm is provided in [43]. Note that sampling under the constraint is similar to directed loop algorithms, which are used to explore the extensive discrete set of ground states in spin models supporting an emergent gauge theory [53], and more efficient algorithms may exist.
The long-time state of the above Glauber dynamics is imposed by the steady-state condition
$\Big{\{}\rho(\theta_{i}+\delta,\varphi_{i}+\delta)w(\theta_{i}+\delta\rightarrow\theta_{i})$
$-\rho(\theta_{i},\varphi_{i})w(\theta_{i}\rightarrow\theta_{i}+\delta)\Big{\}}\Big{|}_{\text{constr.}}=0$ (7)
where $\delta{=}\theta^{\prime}_{i}{-}\theta_{i}$ denotes the small change of the angle $\theta_{i}$ in the Glauber update. We note that detailed balance here is explicitly broken when restoring the constraint, see SM [43]. Using the rates from Eq. (6) and Taylor-expanding to linear order in $\delta$, the density $\rho$ obeys
$\left(\partial_{\varphi_{i}}{+}\partial_{\theta_{i}}\right)\left[\partial_{\theta_{i}}H|_{\text{constr.}}{+}T\left(\partial_{\varphi_{i}}{+}\partial_{\theta_{i}}\right)\right]\rho\left(\theta_{i},\varphi_{i}\right){=}0.$ (8)
We now argue that Eq. (8) arises independently as the Fokker-Planck equation of the original Langevin dynamics at steady state. Recall that the equation of motion (1) is generated by the Hamiltonian (2) under the constraint (5) for each pair $(\theta_{i},\varphi_{i})$. In addition, assume that the two spins $\mathbf{S}_{i}$ and $\mathbf{a}_{i}$ are coupled to the same thermal bath and therefore experience identical noise. Let us rewrite the overdamped Langevin dynamics (1) as $\dot{\alpha}_{i}\left(t\right){=}{-}\left\{\alpha_{i},H\right\}+\sqrt{2T}\eta_{i}(t)$ under the constraint, for $\alpha{=}\theta,\varphi$. One way to recover the configuration-space distribution is by introducing a test function $f(\{\alpha_{j}(t)\})$ and calculating its average value under an ensemble of trajectories generated by the above Langevin equation. In a steady state, the ensemble-average of $f$ is time-independent, i.e., $\partial_{t}\left\langle f(\{\alpha_{j}(t)\})\right\rangle{=}0$. Denoting the corresponding phase-space probability density by $\rho\left(\theta_{i},\varphi_{i}\right)$, this time-translation invariance leads to the textbook steady-state Fokker-Planck equation when the constraint is ignored [54, 55]. However, taking into account the constraint (5) imprints additional correlations between $\theta_{j}$ and $\varphi_{j}$, leading to Eq. (8), see SI [43]. Thus, Glauber Monte-Carlo dynamics generated by the Hamiltonian (2) produces the same steady-state distribution as the Langevin equation (1).
Figure 3: (a) Average magnetization $\langle m\rangle$ and (b) specific heat $C_{v}$ as functions of temperature $T$, obtained from Monte-Carlo simulations with Glauber dynamics based on either the selfish energy (red) or the constrained Hamiltonian embedding (blue), as well as simulations of the original Langevin dynamics (green). Whereas the constrained Glauber and Langevin simulations show the same critical temperature within the statistical uncertainty: $T_{c}=0.78\pm 0.02J$ (Langevin) and $T_{c}=0.79\pm 0.03J$ (constrained Glauber), the selfish-energy Glauber dynamics predicts $T_{c}=0.86\pm 0.02J$. The square lattice has linear dimension $L{=}100$. We use a vision cone angle $\Psi{=}0.85\pi$ and set the time step $\delta t=0.1$ for the Langevin simulation. We use $20L^{4}$ ($3\times 10^{5}$) timesteps for the constrained Glauber (Langevin) dynamics.
We provide evidence for the equivalence of the Langevin and Glauber descriptions in numerical simulations. In Fig. 3, we compare (a) the average magnetization $\langle m\rangle{=}\sqrt{(\sum\cos\theta_{i})^{2}{+}\left(\sum\sin\theta_{i}\right)^{2}}/L^{2}$ and (b) the specific heat, $C_{v}{=}\left(\left\langle H_{SS}^{2}\right\rangle{-}\left\langle H_{SS}\right\rangle^{2}\right)/\left(LT\right)^{2}$, obtained as the variance of the energy in the original spin layer [56], as a function of temperature, for constrained Glauber dynamics (blue) vs. Langevin dynamics (green). We initialize the system in an ordered state, $\theta_{i}{=}0,\varphi_{i}{=}\pi$, with initial energy of the original system $H_{SS}{=}3L^{2}$. Consistently with previous works [39, 29, 40, 41], our data indicates the existence of a phase transition between an ordered phase at low temperature and disordered phase at high temperature [Fig. 3(a)]. This critical behavior is corroborated by the specific heat $C_{v}$, which exhibits a peak at the phase transition point [see Fig. 3(b)]. Importantly, both the Langevin and the constrained Monte-Carlo simulations exhibit the same critical temperature, $T_{c}/J=0.78\pm 0.02$ (Langevin) and $T_{c}/J=0.79\pm 0.03$ (constrained Glauber), for the phase transition, within statistical uncertainty, see SI [43]. By contrast, a Monte-Carlo simulation using the Hamiltonian (3) without including the auxiliary spins $\mathbf{a}_{i}$ produces a critical temperature of $T_{c}/J{\approx}1.72$ [43, 57] (not shown).
Unlike Gauber dynamics based on the selfish energy [30], our constrained-Hamiltonian Monte-Carlo update rule results in the same configuration-space density for the steady state as the Langevin equation. Indeed, the critical temperatures for the phase transition estimated from the Langevin and the selfish-energy approach do not agree in general [compare green and red curves in Fig. 3], although for some parameters the values may agree within statistical uncertainty. By ensuring the equivalence of the Langevin and the constrained Monte-Carlo dynamics, our Hamiltonian embedding resolves this long-standing issue in the study of vision-cone interactions [45].
We emphasize that, whereas we used a specific model to demonstrate the equivalence of the Monte-Carlo dynamics with the original nonreciprocal Langevin equation, the concept is generally applicable to other nonreciprocal systems, as we show in [43].
V Hamiltonian engineering for nonreciprocal systems
The Hamiltonian framework for nonreciprocal interactions that we propose has a natural symplectic structure which, besides applications in Monte-Carlo simulations, enables the use of canonical transformations to study the physics of nonreciprocal systems. As a concrete example, here we demonstrate the application of Floquet engineering to modify the properties of nonreciprocal systems using high-frequency periodic driving [42, 58, 59]. This enables one to reach regimes of interest that might be inaccessible in experiments due to the limited tunability of parameters.
Consider again the vision-cone spins on a square lattice and assume that one has no control over the magnitude and sign of the interaction strength; such situations are abundant in experimental systems. We will now show how high-frequency periodic driving allows us to effectively modify the magnitude and sign of the interaction strength. Let us subject the system to periodic driving, motivated by a condensed-matter setting  [^2] The XY model arises naturally in the description of Josephson junction arrays. When subject to an external AC magnetic field, the latter can be modeled by a time-dependent vector potential which couples to the phase, $\theta_{i}$, of the superconducting order parameter [77]., following the equations of motion
$\dot{\theta}_{i}\left(t\right){=}{-}\!\!\sum_{j\in\{\langle ij\rangle\}}\!\!\!\!J_{ij}(\theta_{i})\sin(\theta_{i}-\theta_{j}-A_{ij}\cos\omega t)+\sqrt{2T}\eta_{i}(t).$ (9)
We model the external periodic drive along the $x$-direction with frequency $\omega$ and amplitude $A_{ij}{=}aA_{0}\cos\psi_{ij}$, where the angle $\psi_{ij}$ is defined in Fig. 2a, and $a$ is the lattice constant [61, 62, 63]. We will show that the periodic drive allows us to effectively suppress the spin interactions along the $x$-direction only [Fig. 4a] while keeping the vision-cone interaction along the $y$-direction — an instance of Floquet engineering known as dynamical localization [42, 58, 59]. In systems with limited control over the interaction strength, this strategy would allow experiments to study the fate of the phase transition, as the dimensionality of the system is effectively changed from $d=2$ (a square lattice) to $d=1$ (a collection of uncoupled chains).
The Hamiltonian formalism proves convenient to analyze the driven system due to its symplectic structure. Adding the auxiliary spin $\mathbf{a}_{j}$, we construct the time-periodic Hamiltonian embedding
$H\left(t\right)=-\sum_{\langle ij\rangle}$ $J_{ij}(\theta_{i})[\cos(\theta_{i}-\theta_{j}-A_{ij}\cos\omega t)$
$+\cos(\varphi_{i}-\theta_{j}-A_{ij}\cos\omega t)]+(i\leftrightarrow j),$
with $\left\{\theta_{j},\varphi_{i}\right\}{=}\delta_{ij}$. To derive an effective description at high driving frequencies, we apply the Floquet-Magnus expansion $H_{F}{=}H_{F}^{(0)}{+}\mathcal{O}(\omega^{-1})$ [42]; the leading-order Floquet Hamiltonian is the time-average
[…]
$+(i\leftrightarrow j),$
where $\mathcal{J}_{0}\left(A_{ij}\right)$ is the Bessel function of the first kind.
Figure 4: The Hamiltonian embedding enables the use of Floquet techniques to analyze driven nonreciprocal interacting systems. (a) Schematic illustration of the periodically driven XY spins (left panel, Eq. (V)) and the effective time-averaged interactions (right panel, Eq. (11)). Thin black dashed lines indicate suppression of interactions in the $x$-direction. (b) Average magnetization $\langle m\rangle$ obtained from Langevin dynamics as a function of temperature $T$ for different drive amplitude strengths $A_{0}$. As $A_{0}$ increases, the transition temperature decreases; as predicted by the time-averaged Floquet Hamiltonian, when $A_{0}{=}z_{n}a^{-1}$ (see text), the system no longer exhibits a long-range ordered phase. We use $10^{5}$ iterations with a timestep of $\delta t=0.01$; $\omega{=}20\pi J$, and $\Psi=7\pi/8$. (c-d) Average magnetization $\langle m\rangle$ obtained from Langevin dynamics as a function of time $t$ for two drive frequencies: $\omega{=}\pi J$ (c) and $\omega{=}20\pi J$ (d), for $T{=}0.5J$. As the frequency increases, the dynamics corresponding to the equations of motion derived from $H_{F}^{(0)}$ (orange) become closer to the exact periodically driven Langevin dynamics (purple). Both show a significant deviation from the Langevin dynamics of the non-driven system from Eq. (1) (green). Parameters: $A_{0}{=}2a^{-1}$, and $L{=}64$.
To verify the validity of the Floquet Hamiltonian description, we study the average magnetization as a function of time $t$ for two driving frequencies in Fig. 4(c,d). The deviation between the results from the exact Langevin equations of motion (purple) and the effective equations of motion derived from $H_{F}^{(0)}$ (red) decreases as the frequency increases [from Fig. 4(c) to Fig. 4(d)]; moreover, the drive modifies the behavior significantly (compare red and green curves).
By tuning the amplitude $A_{0}$ of the drive close to the zero of the Bessel function, $\mathcal{J}_{0}(z_{n}){=}0$, one can suppress spin interactions in the $x$-direction, thereby realizing a dimensional crossover from a two-dimensional lattice to a set of one-dimensional chains [Fig. 4a]. Figure 4(b) shows the average magnetization obtained from simulations of the spin dynamics (9) in the high-frequency regime as a function of temperature $T$. As we gradually increase $A_{0}$, the transition temperature $T_{c}$ decreases. To leading order in $\omega^{-1}$, when $A_{0}{=}z_{n}a^{-1}$, the system decouples into a set of independent 1D XY-chains and no longer exhibits an ordered phase at any finite reservoir temperature. The behavior in Fig. 4(b) is consistent with our analysis using the Floquet Hamiltonian $H_{F}^{(0)}$, since the dimensional crossover occurs precisely at the zeroes of the Bessel function. Finally, note that, even though periodic drives typically lead to enhanced energy absorption, the heating rates are either power-law or exponentially suppressed with increasing the drive frequency $\omega$ [64].
VI Discussion and Outlook
In the past, isolated examples of Hamiltonian descriptions of nonreciprocal interactions have been reported, e.g. for the Lotka-Volterra predator-prey model [65]. In this work, we have constructed a general Hamiltonian embedding for nonreciprocal systems by introducing auxiliary degrees of freedom that couple to the original ones in a reciprocal way. By imposing a constraint on the phase-space dynamics of the embedding, we preserve the total number of independent degrees of freedom and recover the original nonreciprocal interactions. As a direct consequence of our construction, it follows that the phase space of the reciprocal Hamiltonian $H$ from Eq. (2) contains a nonreciprocal submanifold defined by the constraint, which is closed under the dynamics.
According to Liouville’s theorem, without constraint, the Hamiltonian embedding gives rise to nondissipative dynamics in the embedding phase space, even if the original equations of motion are dissipative. This feature is brought about by the symplectic structure inherent to the Hamiltonian embedding (e.g., in the vision-cone XY spins that we discussed, $\theta$ and $\varphi$ are conjugate variables). By contrast, on the submanifold defined by the constraint, the dynamics are dissipative. This apparent contradiction is resolved by noticing that the constrained submanifold is a set of measure zero in the embedding phase space which does not contribute to phase-space volume. However, the mechanism by which dissipation emerges from constrained Hamiltonian dynamics presents an intriguing direction for future research, closely related to a formulation of Hamiltonian dynamics for nonconservative systems [56].
When evaluated on the constraint (5), the total Hamiltonian $H$ vanishes identically; therefore, $H$ does not describe the energy of the original nonreciprocal system, but merely provides a generator for its dynamics. Moreover, the ground-state manifold of $H$ need not obey the constraint. These unusual features arise as a result of the original nonreciprocal system being intrinsically out of equilibrium.
Crucially, our construction works for pairwise interactions, and may not apply to arbitrary non-potential systems [cf. Fig. 1]; a generalization to velocity-dependent (e.g., Lorentz-like) forces is an interesting direction for future work. Unlike in systems where non-reciprocity arises as a result of projecting/tracing out bath degrees of freedom (e.g., a chemical concentration field) that mediate the interactions [1], we emphasize that here the auxiliary degree of freedom does not play the role of a bath for the original one; instead, the projecting/tracing out procedure is replaced by the constraint. The framework is generic and applies to both inertial and dissipative dynamics featuring nonreciprocal pairwise interactions.
We discussed two applications using XY spins on a square lattice with vision-cone interactions: (i) We proved the equivalence between Glauber dynamics based on the Hamiltonian embedding and the original nonreciprocal Langevin dynamics. In particular, we showed that both procedures yield the same critical temperature for the transition between the ferromagnetic and non-magnetized phases. (ii) We make use of the inherent symplectic structure of the embedding to derive an effective Hamiltonian for periodically driven vision-cone XY spins using Floquet theory; this allows us to interpret the disappearance of ferromagnetic order when varying the drive amplitude as a dimensional crossover. These results suggest exciting possibilities for experiments to apply Floquet engineering techniques [42, 58, 59] to investigate regimes in parameter space that are otherwise inaccessible in the nonreciprocal system.
We note that the Hamiltonian embedding does not introduce new physics content per se beyond what is already captured by the equations of motion defining the original nonreciprocal system. However, it enables the study of nonreciprocal systems using analytical and numerical tools from equilibrium or non-equilibrium physics that require a Hamiltonian structure. For instance, one can now use Monte Carlo algorithms that shortcut long transients of Langevin simulations and lead directly to the steady state. Likewise, one can apply various canonical transformations (e.g., Schrieffer-Wolff-like transformations such as the Birkhoff-Gustavson normal form [34, 35, 36]) to derive nonreciprocal models with effective degrees of freedom. An illustrative example is the Floquet-Magnus expansion for periodically driven systems: higher-order terms in the inverse drive frequency contain effective spin-spin interactions [42]. Note that whereas it is possible to derive effective stroboscopic equations of motion in the high-frequency regime directly from the equations of motion to any order in the inverse frequency $\omega^{-1}$ by using two-times perturbation theory [66], the Hamiltonian formulation has the advantage of providing a systematic model-agnostic algorithmic procedure [64]. This formalism can prove useful to analyze the behavior of time-crystalline nonreciprocal states [67].
The Hamiltonian framework we propose has the potential to expand the scope of research on nonreciprocal systems and enable novel applications. In experiments that simulate classical and quantum models [23, 68], our construction offers a new way to engineer nonreciprocal dynamics using reciprocal Hamiltonian systems subject to a constraint. Promising future directions include developing kinetic Monte Carlo schemes to capture non-steady states [38], canonically quantizing the Hamiltonian embedding, or using it to construct statistical mechanics and thermodynamic descriptions of nonreciprocal systems.
Acknowledgments.—We thank L. Hahn, A. McRoberts, M. Sonner, Y. Zhang, and H. Zhao for related discussions. MB was funded by the European Union (ERC, QuSimCtrl, 101113633). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This work was in part supported by the Deutsche Forschungsgemeinschaft via Research Unit FOR 5522 (project-id 499180199) and the cluster of excellence ct.qmat (EXC 2147, Project No. 390858490).
Author contributions.—R.A., M.B., and R.M. conceived the research. Y.S. did the analytical calculations and performed the numerical simulations. R.A. and M.B. supervised the work on the project, and contributed equally to this work. All authors contributed to writing the manuscript.
Appendix A Constrained Hamiltonian embedding of nonreciprocal systems
In this section, we first outline a general approach for constructing Hamiltonian embeddings. We then give the explicit form of the constrained Hamiltonian for various models with nonreciprocal equations of motion.
We begin by emphasizing that, in its simplest form, (non-)reciprocity is a property defined for pairwise interacting systems. Let us denote by ${\bf F}_{i\rightarrow j}(\mathbf{r}_{i},\mathbf{r}_{j})$ the force exerted by particle $i$ at position $\mathbf{r}_{i}$ on particle $j$ at position $\mathbf{r}_{j}$. We consider translation invariant systems where ${\bf F}_{i\rightarrow j}(\mathbf{r}_{i},\mathbf{r}_{j})={\bf F}_{i\rightarrow j}(\mathbf{r}_{i}-\mathbf{r}_{j})$ depends only on the distance between the two particles involved. In addition, we only consider momentum-independent forces, and thus leave out of consideration Lorentz-like forces.
Reciprocal (R) interactions obey Newton’s third law
${\bf F}^{\text{R}}_{i\rightarrow j}(\mathbf{r}_{i}-\mathbf{r}_{j})=-{\bf F}^{\text{R}}_{j\rightarrow i}(\mathbf{r}_{j}-\mathbf{r}_{i}),$ (12)
i.e, the force exerted by particle $i$ on particle $j$, ${\bf F}^{\text{R}}_{i\rightarrow j}(\mathbf{r}_{i}-\mathbf{r}_{j})$, is equal in magnitude and opposite in direction to the force ${\bf F^{\text{R}}}_{j\rightarrow i}(\mathbf{r}_{j}-\mathbf{r}_{i})$ exerted by particle $j$ on particle $i$.
We define nonreciprocal interactions via the relation
${\bf F}^{\text{NR}}_{i\rightarrow j}(\mathbf{r}_{i}-\mathbf{r}_{j})\neq-{\bf F}^{\text{NR}}_{j\rightarrow i}(\mathbf{r}_{j}-\mathbf{r}_{i}).$ (13)
Nonreciprocal interactions refer to an asymmetry in the interaction between two particles in which the action and reaction are not equal.
In general, the equations of motion for particle $i$ may be governed by a combination of reciprocal and nonreciprocal interactions. Hence in its most general form, it takes the form:
$m\ddot{\mathbf{r}}_{i}+\gamma\dot{\mathbf{r}}_{i}=\sum_{j\neq i}\left[\mathbf{F}_{j\rightarrow i}^{\mathrm{NR}}(\mathbf{r}_{i}-\mathbf{r}_{j})+\mathbf{F}_{j\rightarrow i}^{\mathrm{R}}(\mathbf{r}_{i}-\mathbf{r}_{j})\right],$ (14)
where we denote the mass by $m$ and the damping factor by $\gamma$.
We now assume that the interaction forces are conservative, i.e., they can be derived from an interaction potential:
${\bf F}_{i\rightarrow j}(\mathbf{r}_{i}-\mathbf{r}_{j})=-\frac{\partial}{\partial\mathbf{r}_{i}}U_{i\rightarrow j}(\mathbf{r}_{i}-\mathbf{r}_{j}).$ (15)
Reciprocal interactions share the same (mutual) interaction potential, and hence obey the relation $U^{\text{R}}_{i\rightarrow j}=U^{\text{R}}_{j\rightarrow i}$ which, therefore, implies Newton’s third law (Eq. (12)); it is common to use the simplified notation $U^{\text{R}}_{i\rightarrow j}=U^{\text{R}}$. By contrast, nonreciprocal interactions obey $U^{\text{NR}}_{i\rightarrow j}\neq U^{\text{NR}}_{j\rightarrow i}$. Since each interaction function $U^{\text{NR}}_{i\rightarrow j}(\mathbf{r}_{i}-\mathbf{r}_{j})$, $U^{\text{NR}}_{j\rightarrow i}(\mathbf{r}_{i}-\mathbf{r}_{j})$ depends on the positions of both particles involved, one cannot write down an energy function such that the Hamiltonian equations of motion for the corresponding conjugate variables are nonreciprocal; this is because the resulting force acting on particle $i$ will inevitably contain the gradients of both interaction potentials which would make the net force reciprocal.
To resolve this problem, we introduce a copy of the system described by the auxiliary degrees of freedom $\{\mathbf{x}_{i},\mathbf{p}^{\mathbf{x}}_{j}\}=\delta_{ij}$. The original degrees of freedom obey the Poisson bracket $\{\mathbf{r}_{i},\mathbf{p}^{\mathbf{r}}_{j}\}=\delta_{ij}$. One can construct a general Hamiltonian embedding for the joint system using only reciprocal interactions, as follows:
$H$ $=$ $\phantom{+}\sum_{j}\frac{\left(\mathbf{p}_{j}^{\mathbf{r}}-\frac{\gamma}{2}\mathbf{x}_{j}\right)^{2}}{2m}-\frac{\left(\mathbf{p}_{j}^{\mathbf{x}}+\frac{\gamma}{2}\mathbf{r}_{j}\right)^{2}}{2m}$ (16)
$+\sum_{\langle ij\rangle}\left[U_{i\rightarrow j}^{\text{NR}}(\mathbf{r}_{i}-\mathbf{r}_{j})+U_{j\rightarrow i}^{\text{NR}}(\mathbf{r}_{j}-\mathbf{r}_{i})-U_{i\rightarrow j}^{\text{NR}}(\mathbf{x}_{j}-\mathbf{r}_{i})-U_{j\rightarrow i}^{\text{NR}}(\mathbf{x}_{i}-\mathbf{r}_{j})\right]$
$+\sum_{\langle ij\rangle}\left[U^{\text{R}}(\mathbf{r}_{i}-\mathbf{r}_{j})-U^{\text{R}}(\mathbf{x}_{i}-\mathbf{x}_{j})\right].$
The notation $\langle ij\rangle$ denotes a summation over pairs of particles $i$ and $j$ (each counted once). Here the kinetic energy is quadratic in the momenta $\mathbf{p}_{j}^{\mathbf{r}},\mathbf{p}_{j}^{\mathbf{x}}$, and hence the system is inertial and the EOM is second order in time. Note the apparent similarity of the kinetic energy above to that of a particle in a magnetic field; however, a notable difference is that in the Hamiltonian $H$ above the minimal coupling is between $(\mathbf{p}_{j}^{\mathbf{r}},\mathbf{x}_{j})$ [as opposed to $(\mathbf{p}_{j}^{\mathbf{r}},\mathbf{r}_{j})$], etc.
As we showed in the Main Text, in general, the equations of motion derived from the Hamiltonian $H$ defines a coupled system of equations
$\frac{\mathrm{d}\mathbf{r}_{i}}{\mathrm{d}t}$ $=$ $\frac{\mathbf{p}_{i}^{\mathbf{r}}-\frac{\gamma}{2}\mathbf{x}_{i}}{m}\ ,$
$\frac{\mathrm{d}\mathbf{p}_{i}^{\mathbf{r}}}{\mathrm{d}t}$ $=$ $\frac{\gamma}{2}\frac{\left(\mathbf{p}_{i}^{\mathbf{x}}+\frac{\gamma}{2}\mathbf{r}_{i}\right)}{m}-\frac{\partial}{\partial{\mathbf{r}_{i}}}\sum_{j\in\langle ij\rangle}\left[U_{i\rightarrow j}^{\text{NR}}(\mathbf{r}_{i}-\mathbf{r}_{j})+U_{j\rightarrow i}^{\text{NR}}(\mathbf{r}_{j}-\mathbf{r}_{i})-U_{i\rightarrow j}^{\text{NR}}(\mathbf{x}_{j}-\mathbf{r}_{i})+U^{\text{R}}(\mathbf{r}_{i}-\mathbf{r}_{j})\right]\ ,$
$\frac{\mathrm{d}\mathbf{x}_{i}}{\mathrm{d}t}$ $=$ $-\frac{\mathbf{p}_{i}^{\mathbf{x}}+\frac{\gamma}{2}\mathbf{r}_{i}}{m}\ ,$
$\frac{\mathrm{d}\mathbf{p}_{i}^{\mathbf{x}}}{\mathrm{d}t}$ $=$ $\frac{\gamma}{2}\frac{\left(\mathbf{p}_{i}^{\mathbf{r}}-\frac{\gamma}{2}\mathbf{x}_{i}\right)}{m}+\frac{\partial}{\partial{\mathbf{x}_{i}}}\sum_{j\in\langle ij\rangle}\left[U_{j\rightarrow i}^{\text{NR}}(\mathbf{x}_{i}-\mathbf{r}_{j})+U^{\text{R}}(\mathbf{x}_{i}-\mathbf{x}_{j})\right].$ (17)
However, if we impose the constraint
$\mathbf{r}_{i}(0)-\mathbf{x}_{i}(0)=0,\qquad\mathbf{p}_{i}^{\mathbf{r}}(0)+\mathbf{p}_{i}^{\mathbf{x}}(0)=0,$ (18)
at time $t=0$, it follows immediately from the above equations of motion that the constraint is preserved at all times, i.e.,
$\frac{\mathrm{d}}{\mathrm{d}t}\left[{\mathbf{r}}_{i}(t)-{\mathbf{x}}_{i}(t)\right]\bigg{|}_{\text{constr.}}=0,\qquad\frac{\mathrm{d}}{\mathrm{d}t}\left[{\mathbf{p}}_{i}^{\mathbf{r}}(t)+{\mathbf{p}}_{i}^{\mathbf{x}}(t)\right]\bigg{|}_{\text{constr.}}=0.$ (19)
Moreover, the joint equations of motion reduce to the original nonreciprocal dynamics in Eq. (14). More precisely, each degree of freedom (i.e., both the $\mathbf{r}_{j}$ and the $\mathbf{x}_{j}$ degree of freedom) follows the same equations of motion; the dynamics of the $\mathbf{x}_{j}$ degree of freedom perfectly mirrors that of $\mathbf{r}_{j}$. We note that the negative kinetic energy for the auxiliary $\mathbf{x}_{j}$ is a sufficient condition to preserve the constraint at all times.
Below, we discuss concrete examples of this general Hamiltonian embedding for nonreciprocal interactions, see Fig. 5.
Figure 5: Schematic representation of various models exhibiting nonreciprocal equations of motion. (a) Metal-dielectric self-propelled Janus colloidal particles (Sec. A.2) experience self-propulsion, reciprocal repulsive forces and nonreciprocal torques. (b) Self-aligning active particles (Sec. A.3) interact via reciprocal repulsive forces and exhibit self-propulsion. The nonreciprocal torques cause each particle to rotate toward its velocity direction. (c) Programmed robotic metamaterials (Sec. A.4) consist of robotic units connected by springs that exert asymmetric forces on the left and right sides. (d) Feedback-controlled active particles (Sec. A.5) interact via reciprocal forces and self-propel when a sufficient number of neighboring particles enter their vision cone.
A.1 XY spins with vision cone interactions
A first illustrative example is the vision-cone XY spins, $\mathbf{S}_{i}=(\cos\theta_{i},\sin\theta_{i})$, on a square lattice discussed in the Main Text. Here, the nonreciprocal equations of motion take the form [cf. Eq. (1)]:
$\dot{\theta_{i}}=-\sum_{j\in\langle ij\rangle}J_{ij}(\theta_{i})\sin(\theta_{i}-\theta_{j})+\sqrt{2T}\eta_{i}.$ (20)
To construct the Hamiltonian, we introduce the auxiliary spins $\mathbf{a}_{i}$ parametrized by the angle $\varphi_{i}$. Proceeding as explained in the Main Text, we obtain:
$H$ $=$ $-\sum_{\langle ij\rangle}\left[J_{ij}(\theta_{i})(\mathbf{S}_{i}\cdot\mathbf{S}_{j}+\mathbf{a}_{i}\cdot\mathbf{S}_{j})+J_{ji}(\theta_{j})(\mathbf{S}_{j}\cdot\mathbf{S}_{i}+\mathbf{a}_{j}\cdot\mathbf{S}_{i})\right],$ (21)
$=$ $-\sum_{\langle ij\rangle}\left\{J_{ij}(\theta_{i})[\cos(\theta_{i}-\theta_{j})+\cos(\varphi_{i}-\theta_{j})]+J_{ji}(\theta_{j})[\cos(\theta_{j}-\theta_{i})+\cos(\varphi_{j}-\theta_{i})]\right\}.$
The equations of motion are given in Eq. (4), ignoring the noise term. The origin and technical handling of the constraint are discussed in detail in Sec. B.
As we now demonstrate, this way of constructing the Hamiltonian is generic and can be applied to other models with nonreciprocal interactions.
A.2 Metal-dielectric Janus particles
The next example is that of metal-dielectric self-propelled Janus colloidal particles. These particles are silica spheres coated with titanium on one hemisphere. Suspended in water, they sediment and form a monolayer on a conductive coverslip, separated from another conductive coverslip, which together form a capacitor [69, 15, 20]. The particles are driven by an AC electric field produced by a voltage between the conductive coverslips. The electric field induces dipoles in each of the particles’ hemispheres, which leads both to self-propulsion as well as to electrostatic interactions between the particles. The interactions consist of a repulsive force and a torque that tends to make the particles turn either towards or away from one another [15, 20]. At low area fractions and self-propulsion speeds, the system behaves as an isotropic active gas. At higher area fractions and speeds, the system either undergoes phase separation when particles turn towards each other [15] or transitions to a polar-ordered flock when particles turn away from one another [20].
To explain these phenomena, previous work proposed a two-dimensional microscopic model based on the dipolar interactions between the hemispheres of the particles (Fig. 5a) [15]. The interaction torques, which tend to rotate a particle either away or towards the position of a second particle, are intrinsically nonreciprocal; they do not have a reciprocal limit [15, 20]. The position and orientation of particle $i$ are denoted by $\mathbf{r}_{i}$ and $\mathbf{\hat{n}}_{i}$, respectively. The equations of motion are given by
$\frac{\mathrm{d}\mathbf{r}_{i}}{\mathrm{d}t}$ $=$ $v_{0}\mathbf{\hat{n}}_{i}+\sum_{j\neq i}\frac{\mathbf{F}_{ji}}{\gamma_{\text{t}}}+\bm{\eta}_{i}^{\text{t}},$
$\frac{\mathrm{d}\mathbf{\hat{n}}_{i}}{\mathrm{d}t}$ $=$ $\sum_{j\neq i}\frac{\mathbf{\Gamma}_{ji}}{\gamma_{\text{r}}}+\hat{\mathbf{n}}_{i}\times\bm{\eta}_{i}^{\text{r}}.$ (22)
Here, $\bm{\eta}_{i}^{\text{t}}$ and $\bm{\eta}_{i}^{\text{r}}$ are translational and rotational Gaussian white noises with zero mean and with $\langle\eta^{\text{t}}_{i,\alpha}(t)\eta^{\text{t}}_{j,\beta}(t^{\prime})\rangle=2D_{\text{t}}\delta_{ij}\delta_{\alpha\beta}\delta(t-t^{\prime})$ and $\langle\eta^{\text{r}}_{i,z}(t)\eta^{\text{r}}_{j,z}(t^{\prime})\rangle=2D_{\text{r}}\delta_{ij}\delta(t-t^{\prime})$, where $D_{\text{t}}$ and $D_{\text{r}}$ are the translational and rotational diffusion coefficients of the particles. Here, Latin indices indicate particles, whereas Greek indices indicate spatial components. Note that the rotational noise has components only along the $\hat{\mathbf{z}}$ axis so that the particle orientation $\hat{\mathbf{n}}_{i}$ remains on the $xy$ plane.
The particle positions change due to self-propulsion at a speed $v_{0}$, and due to the reciprocal repulsive force $\mathbf{F}_{ji}$ exerted by particle $j$ on particle $i$, with $\gamma_{\text{t}}$ being the translational friction coefficient. This interaction force can be derived from an interaction potential $U({|\mathbf{r}_{i}-\mathbf{r}_{j}|})$ between particles $i$ and $j$. In addition, $\mathbf{\Gamma}_{ji}$ denotes the interaction torque, which is intrinsically nonreciprocal, $\mathbf{\Gamma}_{ji}\neq-\mathbf{\Gamma}_{ij}$, and $\gamma_{\text{r}}$ is the associated rotational friction coefficient.
Assuming turn-towards torques, the expressions for the force $\mathbf{F}_{ji}$ and the torque $\mathbf{\Gamma}_{ji}$ are given by [15]
$\mathbf{F}_{ji}$ $=$ $\frac{3}{4\pi\epsilon}\frac{\left(d_{\text{h}}+d_{\text{t}}\right)^{2}}{r_{ji}^{4}}\exp\left(-\frac{r_{ji}}{\lambda}\right)\hat{\mathbf{r}}_{ji},$
$\mathbf{\Gamma}_{ji}$ $=$ $\frac{3\ell}{4\pi\epsilon}\frac{d_{\text{h}}^{2}-d_{\text{t}}^{2}}{r_{ji}^{4}}\exp\left(-\frac{r_{ji}}{\lambda}\right)\mathbf{\hat{n}}_{i}\times\hat{\mathbf{r}}_{ji}=\bm{\tau}_{ji}\times\mathbf{\hat{n}}_{i},$ (23)
where $\epsilon$ is the dielectric permittivity of the solvent, and $d_{\text{h}}$ and $d_{\text{t}}$ are the effective dipole strengths of the head and tail hemispheres. Here, to shorten the notation, we have rewritten the interaction torque in terms of the quantity
$\bm{\tau}_{ji}\equiv-\frac{3\ell}{4\pi\epsilon}\frac{d_{\text{h}}^{2}-d_{\text{t}}^{2}}{r_{ji}^{4}}\exp\left(-\frac{r_{ji}}{\lambda}\right)\hat{\mathbf{r}}_{ji}.$ (24)
Here $\mathbf{r}_{ji}=\mathbf{r}_{i}-\mathbf{r}_{j}$ is the distance vector with its length and direction given by $r_{ji}=\left|\mathbf{r}_{ji}\right|$ and $\hat{\mathbf{r}}_{ji}=\mathbf{r}_{ji}/r_{ji}$, respectively.
So far, we discussed all forces and torques. We now introduce the auxiliary particles with position and orientation denoted as $\mathbf{x}_{i}$ and $\mathbf{\hat{m}}_{i}$. The Hamiltonian embedding is given by
$H$ $=$ $\sum_{i}-v_{0}\mathbf{\hat{n}}_{i}\cdot\left(\mathbf{r}_{i}+\mathbf{x}_{i}\right)+\sum_{i\neq j}\left[\frac{U(|\mathbf{r}_{i}-\mathbf{r}_{j}|)}{{\gamma_{\text{t}}}}-\frac{U(|\mathbf{x}_{i}-\mathbf{x}_{j}|)}{{\gamma_{\text{t}}}}+\frac{\bm{\tau}_{ji}}{\gamma_{\text{r}}}\cdot\left(\mathbf{\hat{n}}_{i}+\mathbf{\hat{m}}_{i}\right)+\frac{\bm{\tau}_{ij}}{\gamma_{\text{r}}}\cdot\left(\mathbf{\hat{n}}_{j}+\mathbf{\hat{m}}_{j}\right)\right].$ (25)
We can derive the corresponding equations of motion from the Hamiltonian description. The translational motion of the particles is governed by
$\frac{\mathrm{d}\mathbf{r}_{i}}{\mathrm{d}t}=-\frac{\partial H}{\partial\mathbf{r}_{i}}=v_{0}\mathbf{\hat{n}}_{i}-\sum_{j\neq i}\left[\frac{1}{{\gamma_{\text{t}}}}\frac{\partial U(|\mathbf{r}_{i}-\mathbf{r}_{j}|)}{\partial\mathbf{r}_{i}}-\frac{\partial}{\partial\mathbf{r}_{i}}\frac{\bm{\tau}_{ij}}{\gamma_{\text{r}}}\cdot\left(\mathbf{\hat{n}}_{i}+\mathbf{\hat{m}}_{i}\right)-\frac{\partial}{\partial\mathbf{r}_{i}}\frac{\bm{\tau}_{ji}}{\gamma_{\text{r}}}\cdot\left(\mathbf{\hat{n}}_{j}+\mathbf{\hat{m}}_{j}\right)\right],$ (26)
and the orientational motion of the particle follows
$\frac{\mathrm{d}\mathbf{\hat{n}}_{i}}{\mathrm{d}t}=\frac{\partial H}{\partial\mathbf{\hat{n}}_{i}}\times\mathbf{\hat{n}}_{i}=-v_{0}\left(\mathbf{r}_{i}+\mathbf{x}_{i}\right)\times\mathbf{\hat{n}}_{i}+\sum_{j\neq i}\frac{\bm{\tau}_{ji}}{\gamma_{\text{r}}}\times\mathbf{\hat{n}}_{i},$ (27)
where we omit the corresponding equations for $\mathbf{x}_{i}$ and $\mathbf{m}_{i}$.
Applying the constraints
$\mathbf{r}_{i}(t)+\mathbf{x}_{i}(t)$ $=$ $0,$
$\mathbf{\hat{n}}_{i}(t)+\mathbf{\hat{m}}_{i}(t)$ $=$ 00 0 (28)
to the equations of motion yields
$\frac{\mathrm{d}\mathbf{r}_{i}}{\mathrm{d}t}$ $=$ $v_{0}\mathbf{\hat{n}}_{i}+\sum_{j\neq i}\frac{\mathbf{F}_{ji}}{\gamma_{\text{t}}},$
$\frac{\mathrm{d}\mathbf{\hat{n}}_{i}}{\mathrm{d}t}$ $=$ $\sum_{j\neq i}\frac{\mathbf{\Gamma}_{ji}}{\gamma_{\text{r}}},$ (29)
which is the same as Eq. (A.2), except for the noise terms $\bm{\eta}_{i}^{\text{t}},\bm{\eta}_{i}^{\text{r}}$.
A.3 Self-aligning active particles
Next we study systems of self-aligning active particles — a model that has been proposed for cells, vibrated disks, and walker robots [70, 71, 44, 46]. The interplay of self-alignment and central forces between the particles has been found to give rise to flocking or collective oscillations, depending on whether the system is confined, and whether it is liquid or solid [70, 71, 72, 44, 46].
Self-aligning active particles, with positions $\mathbf{r}_{i}$ and orientations $\mathbf{\hat{n}}_{i}$, tend to align their orientation towards their velocity. They interact via central forces, arising for example from hard-disk collisions or elastic springs, resulting in nonreciprocal effective interactions. The equations of motion are given by [44, 46]
$\frac{\mathrm{d}\mathbf{r}_{i}}{\mathrm{d}t}$ $=$ $v_{0}\mathbf{\hat{n}}_{i}+\sum_{j\neq i}\frac{\mathbf{F}_{ji}}{\gamma_{\text{t}}}+\bm{\eta}_{i}^{\text{t}},$
$\frac{\mathrm{d}\mathbf{\hat{n}}_{i}}{\mathrm{d}t}$ $=$ $\frac{1}{\ell_{\text{a}}}\left(\mathbf{\hat{n}}_{i}\times\dot{\mathbf{r}}_{i}\right)\times\mathbf{\hat{n}}_{i}+\mathbf{\hat{n}}_{i}\times\bm{\eta}_{i}^{\text{r}}.$ (30)
Here, $v_{0}$ denotes the self-propulsion speed, $\gamma_{\text{t}}$ is the translational friction coefficient, and $\mathbf{F}_{ji}=-\mathbf{F}_{ij}$ represents the reciprocal repulsive force that particle $i$ feels from particle $j$. Such interaction force can be derived from a potential $U(|\mathbf{r}_{i}-\mathbf{r}_{j}|)$ between particles $i$ and $j$. The second equation expresses that the orientation $\hat{\mathbf{n}}_{i}$ of a particle rotates towards its velocity $\dot{\mathbf{r}}_{i}$ over an alignment length $\ell_{\text{a}}$. A schematic illustration is presented in Fig. 5b. Ignoring the noise terms $\bm{\eta}_{i}^{\text{t}}$ and $\bm{\eta}_{i}^{\text{r}}$, we substitute the velocity in the second equation using the first equation, thereby obtaining the following expression for the equations of motion
$\frac{\mathrm{d}\mathbf{\hat{n}}_{i}}{\mathrm{d}t}=\frac{1}{\ell_{\text{a}}\gamma_{\text{t}}}\sum_{j\neq i}\left(\mathbf{\hat{n}}_{i}\times\mathbf{F}_{ji}\right)\times\mathbf{\hat{n}}_{i}.$ (31)
This expression shows that the self-aligning torque tends to align the particle orientation with the interaction force. Such a torque is intrinsically nonreciprocal because, in general, $\left[\mathbf{\hat{n}}_{i}\times\mathbf{F}_{ji}\right]\times\mathbf{\hat{n}}_{i}\neq-\left[\mathbf{\hat{n}}_{j}\times\mathbf{F}_{ij}\right]\times\mathbf{\hat{n}}_{j}$ as $\hat{\mathbf{n}}_{i}$, and $\hat{\mathbf{n}}_{j}$ are in general different.
We now introduce auxiliary particles with position and orientation denoted by $\mathbf{x}_{i}$ and $\mathbf{\hat{m}}_{i}$. The Hamiltonian embedding is given by
$H$ $=$ $-\sum_{i}v_{0}\mathbf{\hat{n}}_{i}\cdot\left(\mathbf{r}_{i}+\mathbf{x}_{i}\right)$ (32)
$+\frac{1}{\gamma_{\text{t}}}\sum_{i\neq j}\left\{U(|\mathbf{r}_{i}-\mathbf{r}_{j}|)-U(|\mathbf{x}_{i}-\mathbf{x}_{j}|)+\frac{1}{\ell_{\text{a}}}\left[\mathbf{\hat{n}}_{i}\times\mathbf{F}_{ji}\right]\cdot\left(\mathbf{\hat{n}}_{i}+\mathbf{\hat{m}}_{i}\right)+\frac{1}{\ell_{\text{a}}}\left[\mathbf{\hat{n}}_{j}\times\mathbf{F}_{ij}\right]\cdot\left(\mathbf{\hat{n}}_{j}+\mathbf{\hat{m}}_{j}\right)\right\}.$
We can now derive the corresponding equations of motion from this Hamiltonian description. The translational motion of the particles obeys
$\frac{\mathrm{d}\mathbf{r}_{i}}{\mathrm{d}t}$ $=$ $-\frac{\partial H}{\partial\mathbf{r}_{i}}$ (33)
$=$ $v_{0}\mathbf{\hat{n}}_{i}+\sum_{j\neq i}\left\{-\frac{1}{\gamma_{\text{t}}}\frac{\partial U(|\mathbf{r}_{i}-\mathbf{r}_{j}|)}{\partial\mathbf{r}_{i}}+\frac{1}{\ell_{\text{a}}\gamma_{\text{t}}}\frac{\partial\mathbf{F}_{ij}}{\partial\mathbf{r}_{i}}\left[\left(\mathbf{\hat{n}}_{j}+\mathbf{\hat{m}}_{j}\right)\times\mathbf{\hat{n}}_{j}\right]+\frac{1}{\ell_{\text{a}}\gamma_{\text{t}}}\frac{\partial\mathbf{F}_{ji}}{\partial\mathbf{r}_{i}}\left[\left(\mathbf{\hat{n}}_{i}+\mathbf{\hat{m}}_{i}\right)\times\mathbf{\hat{n}}_{i}\right]\right\},$
and the orientational motion of the particle follows
$\frac{\mathrm{d}\mathbf{\hat{n}}_{i}}{\mathrm{d}t}$ $=$ $\frac{\partial H}{\partial\mathbf{\hat{n}}_{i}}\times\mathbf{\hat{n}}_{i}$ (34)
$=$ $-v_{0}\left(\mathbf{r}_{i}+\mathbf{x}_{i}\right)\times\mathbf{\hat{n}}_{i}+\frac{1}{\ell_{\text{a}}\gamma_{\text{t}}}\sum_{j\neq i}\left\{\left[\mathbf{\hat{n}}_{i}\times\mathbf{F}_{ji}\right]\times\mathbf{\hat{n}}_{i}+\left[\mathbf{F}_{ji}\times\left(\mathbf{\hat{n}}_{i}+\mathbf{\hat{m}}_{i}\right)\right]\times\mathbf{\hat{n}}_{i}\right\}.$
Finally, the constraints read as
$\mathbf{r}_{i}(t)+\mathbf{x}_{i}(t)=0,$
$\mathbf{\hat{n}}_{i}(t)+\mathbf{\hat{m}}_{i}(t)=0.$ (35)
Applying the constraints to the equations of motion yields
$\frac{\mathrm{d}\mathbf{r}_{i}}{\mathrm{d}t}$ $=$ $v_{0}\mathbf{\hat{n}}_{i}+\frac{1}{\gamma_{\text{t}}}\sum_{j\neq i}\mathbf{F}_{ij},$
$\frac{\mathrm{d}\mathbf{\hat{n}}_{i}}{\mathrm{d}t}$ $=$ $\frac{1}{\ell_{\text{a}}\gamma_{\text{t}}}\sum_{j\neq i}\left(\mathbf{\hat{n}}_{i}\times\mathbf{F}_{ij}\right)\times\mathbf{\hat{n}}_{i},$ (36)
which is equivalent to Eq. (A.3), except for the noise terms.
A.4 Programmed robotic metamaterials
Let us now look at mechanical metamaterials made of robots whose interactions are programmed to break reciprocity  [23, 24]. In such materials, nonreciprocal interactions give rise to spatially asymmetric standing waves and unidirectionally amplified propagating waves, realizing the mechanical analogue of the non-Hermitian skin effect [73, 74]. Specifically, the robots are programmed to turn more (less) depending on how their left (right) neighbors turn, as schematized in Fig. 5c. The dynamics can be mapped onto the following nonreciprocal discrete wave equation, which arises from a mass-spring model in which the forces exerted by the springs on the left and right sides differ [23]:
$\ddot{u}_{i}$ $=$ $-k\left(1+\varepsilon\right)\left(u_{i-1}-u_{i}\right)-k\left(1-\varepsilon\right)\left(u_{i+1}-u_{i}\right).$ (37)
Here $u_{i}$ denotes the displacement of robot $i$; $k\left(1+\varepsilon\right)$ and $k\left(1-\varepsilon\right)$ represent the stiffnesses of coupling springs from left to right and from right to left, respectively, with the dimensionless quantity $\varepsilon$ quantifying non-reciprocity. For convenience, we have set the mass to unity.
We now introduce the conjugate momentum $p_{i}^{u}$ for $u_{i}$, along with the auxiliary degrees of freedom $p_{i}^{w}$ and $w_{i}$ together with the Poisson bracket relations, $\left\{u_{i},p_{j}^{u}\right\}=\delta_{ij}$, and $\left\{w_{i},p_{j}^{w}\right\}=\delta_{ij}$. A Hamiltonian can then be constructed as follows:
$H=\sum_{i}\left[\frac{\left(p_{i}^{u}\right)^{2}}{2}-\frac{\left(p_{i}^{w}\right)^{2}}{2}+k\left(1+\varepsilon\right)\left(u_{i-1}-u_{i}\right)\left(u_{i}+w_{i}\right)+k\left(1-\varepsilon\right)\left(u_{i+1}-u_{i}\right)\left(u_{i}+w_{i}\right)\right].$ (38)
We observe that the kinetic energy of the auxiliary particles is negative. Such a negative kinetic energy term in the auxiliary system is necessary to enforce our constraints; a more detailed explanation will be provided in Sec. B.2.
We can derive the corresponding equations of motion from the Hamiltonian description
$\dot{u}_{i}$ $=$ $\{u_{i},H\}=p_{i}^{u},$
$\dot{w}_{i}$ $=$ $\{w_{i},H\}=-p_{i}^{w},$
$\dot{p}_{i}^{u}$ $=$ $\{p_{i}^{u},H\}=-k\left(1+\varepsilon\right)\left(u_{i-1}-u_{i}\right)-k\left(1-\varepsilon\right)\left(u_{i+1}-u_{i}\right)$
$+k\left(1+\varepsilon\right)\left(u_{i}+w_{i}\right)+k\left(1-\varepsilon\right)\left(u_{i}+w_{i}\right)-k\left(1+\varepsilon\right)\left(u_{i+1}+w_{i+1}\right)-k\left(1-\varepsilon\right)\left(u_{i-1}+w_{i-1}\right),$
$\dot{p}_{i}^{w}$ $=$ $\{p_{i}^{w},H\}=-k\left(1+\varepsilon\right)\left(u_{i-1}-u_{i}\right)-k\left(1-\varepsilon\right)\left(u_{i+1}-u_{i}\right).$ (39)
The constraint here involves the momenta as well, and read as
$u_{i}(t)+w_{i}(t)$ $=$ $0,$
$p_{i}^{u}(t)-p_{i}^{w}(t)$ $=$ $0$ (40)
Applying the constraint to the equations of motion yields
$\ddot{u}_{i}$ $=$ $-k\left(1+\varepsilon\right)\left(u_{i-1}-u_{i}\right)-k\left(1-\varepsilon\right)\left(u_{i+1}-u_{i}\right).$ (41)
The equation of motion of $u_{i}$ is the same as Eq. (37).
A.5 Feedback-controlled active particles with vision-cone interactions
We next consider light-activated active particles whose motility is controlled by an external feedback loop, consisting of a particle detection algorithm and a scanning laser system [21, 22]. Through this feedback system, particles can be programmed to self-propel only when there are enough other particles in their vision cone [22], thus mimicking the behavior of animals such as midges. To decide whether a given particle $i$ should self-propel or not, a perception metric is computed as
$P_{i}=\sum_{j\in V_{i}^{\alpha}}\frac{1}{2\pi r_{ij}},$ (42)
where $V_{i}^{\alpha}$ denotes the vision cone of angle $\alpha$ around the orientation of particle $i$ (Fig. 5d). Respectively, $r_{ij}$ is the distance between particles $i$ and $j$, so that the perception metric assigns less weight to more distant particles. If a given particle has a perception metric $P_{i}>P^{*}$, it is made to propel at speed $v_{0}$, which is achieved by shining light on it through the laser scanning system. Otherwise, the particle remains passive. Thus, particle motion can be modeled as
$\frac{\mathrm{d}\mathbf{r}_{i}}{\mathrm{d}t}$ $=$ $v\left(P_{i}\right)\hat{\mathbf{n}}_{i}-\sum_{j\neq i}\frac{1}{\gamma_{\text{t}}}\frac{\partial U(r_{ij})}{\partial\mathbf{r}_{i}}+\bm{\eta}_{i}^{\text{t}},$
$\frac{\mathrm{d}\mathbf{\hat{n}}_{i}}{\mathrm{d}t}$ $=$ $\hat{\mathbf{n}}_{i}\times\bm{\eta}_{i}^{\text{r}},$ (43)
where the self-propulsion speed $v$ is given by
$v(P_{i})=\left\{\begin{array}[]{ll}v_{0};&P_{i}>P^{*}\\ 0;&P_{i}<P^{*}.\end{array}\right.$ (44)
Respectively, $\gamma_{\text{t}}$ is the translational friction coefficient, $\bm{\eta}_{i}^{\text{t}}$ and $\bm{\eta}_{i}^{\text{r}}$ are the translational and rotational noises, and $U(r_{ij})$ is an interaction potential to model excluded-volume interactions, which can be done through the repulsive Weeks-Chandler-Andersen (WCA) potential given by
$U(r_{ij})=\left\{\begin{array}[]{lr}4\epsilon\left[\left(\frac{\sigma}{r_{ij}}\right)^{12}-\left(\frac{\sigma}{r_{ij}}\right)^{6}+\frac{1}{4}\right],&\text{for }r_{ij}\leq\sigma\sqrt[6]{2}\\ 0,&\ \ \text{for }r_{ij}>\sigma\sqrt[6]{2},\end{array}\right.$ (45)
where $\epsilon$ is the potential strength and $\sigma$ is the particle diameter.
We now introduce auxiliary particles with position and orientation denoted by $\mathbf{x}_{i}$ and $\mathbf{\hat{m}}_{i}$, respectively, to construct the Hamiltonian embedding
$H=-\sum_{i}v\left(P_{i}\right)\hat{\mathbf{n}}_{i}\cdot\left(\mathbf{r}_{i}+\mathbf{x}_{i}\right)+\frac{1}{\gamma_{\text{t}}}\sum_{i<j}\left[U(r_{ij})-U(x_{ij})\right].$ (46)
We can derive the equations of motion for the particle positions and orientations from the Hamiltonian description as follows
$\frac{\mathrm{d}\mathbf{r}_{i}}{\mathrm{d}t}$ $=$ $-\frac{\partial H}{\partial\mathbf{r}_{i}}=v\left(P_{i}\right)\hat{\mathbf{n}}_{i}+\sum_{j}\frac{\partial v\left(P_{j}\right)}{\partial\mathbf{r}_{i}}\hat{\mathbf{n}}_{j}\cdot\left(\mathbf{r}_{j}+\mathbf{x}_{j}\right)-\sum_{j\neq i}\frac{1}{\gamma_{\text{t}}}\frac{\partial U(r_{ij})}{\partial\mathbf{r}_{i}},$
$\frac{\mathrm{d}\mathbf{\hat{n}}_{i}}{\mathrm{d}t}$ $=$ $\frac{\partial H}{\partial\mathbf{\hat{n}}_{i}}\times\mathbf{\hat{n}}_{i}=v\left(P_{i}\right)\left(\mathbf{r}_{i}+\mathbf{x}_{i}\right)\times\mathbf{\hat{n}}_{i}.$ (47)
Finally, the constraints read as
$\mathbf{r}_{i}(t)+\mathbf{x}_{i}(t)$ $=$ $0,$
$\mathbf{\hat{n}}_{i}(t)+\mathbf{\hat{m}}_{i}(t)$ $=$ $0$ (48)
Applying the constraint to the equations of motion yields
$\frac{\mathrm{d}\mathbf{r}_{i}}{\mathrm{d}t}$ $=$ $v\left(P_{i}\right)\hat{\mathbf{n}}_{i}-\frac{1}{\gamma_{\text{t}}}\sum_{j\neq i}\frac{\partial U(r_{ij})}{\partial\mathbf{r}_{i}},$
$\frac{\mathrm{d}\mathbf{\hat{n}}_{i}}{\mathrm{d}t}$ $=$ $0,$ (49)
which is the same as Eq. (A.5), except for the noise terms.
Appendix B Symplectic structure of the Hamiltonian embedding: Implementing the constraint for interactions with overdamped and undamped nonreciprocal dynamics
In Sec. A, we introduced Hamiltonian embeddings and outlined the appropriate constraints for several nonreciprocal systems, both for damped and undamped dynamics. Here, we elaborate on these two cases separately, discussing in detail the implementation of the constraint and the corresponding symplectic structure. Although the discussion below is general, for concreteness we shall illustrate the procedure using the vision-cone XY spins as an example, for which the interaction $J_{ij}(\theta_{i})$ takes the form
$J_{ij}(\theta_{i})=\left\{\begin{array}[]{cc}J,&\text{for}\;\min\left\{2\pi-\left|\theta_{i}-\psi_{ij}\right|,\left|\theta_{i}-\psi_{ij}\right|\right\}\leq\Psi\\ 0,&\mathrm{else.}\end{array}\right.$ (50)
Below, we start by introducing the form of the Lagrangian and then derive the corresponding Hamiltonian, from which the dissipative and inertial nonreciprocal equations of motion can be obtained, along with the associated Poisson bracket relations. Subsequently, we impose the appropriate constraint and demonstrate that it is preserved at all times.
B.1 Overdamped nonreciprocal systems
Consider a dissipative system with nonreciprocal interactions in the equations of motion (20). Since the equation of motion is first-order in $\theta$, we consider a kinetic energy term linear in $\dot{\theta}$, expressed as $\mathcal{K}=\sum_{i}\varphi_{i}\dot{\theta}_{i}$. Such kinetic energy terms are commonly used to describe charged particles on a plane exposed to a perpendicular magnetic field [75]. In addition, we denote by $\mathcal{U}$ the symmetrized interactions, cf. Eq. (21). The explicit form of the Lagrangian is given by
$\mathcal{L}$ $=$ $\mathcal{K}-\mathcal{U}$ (51)
$=$ $\sum_{i}\varphi_{i}\dot{\theta}_{i}+\sum_{\langle ij\rangle}\left\{J_{ij}\left(\theta_{i}\right)\left[\cos(\theta_{i}-\theta_{j})+\cos(\varphi_{i}-\theta_{j})\right]+J_{ji}\left(\theta_{j}\right)\left[\cos(\theta_{j}-\theta_{i})+\cos(\varphi_{j}-\theta_{i})\right]\right\}.$
From this Lagrangian, we can derive the canonical momentum associated with $\theta_{i}$ as
$p_{i}=\frac{\partial\mathcal{L}}{\partial\dot{\theta}_{i}}=\varphi_{i},$ (52)
which leads to the Poisson bracket relation $\left\{\theta_{i},\varphi_{j}\right\}=\delta_{ij}$. This implies that the original spins $\theta_{i}$ and the auxiliary spins $\varphi_{i}$ are conjugate degrees of freedom. Therefore, they define a phase space that is a torus (i.e., compact).
Using the definition for the canonical momenta, we can construct the corresponding Hamiltonian from the Lagrangian $\mathcal{L}$ by performing a Legendre transform
$\mathcal{H}$ $=$ $\sum_{i}p_{i}\dot{\theta}_{i}-\mathcal{L}=\mathcal{U}$ (53)
$=$ $-\sum_{\langle ij\rangle}\left\{J_{ij}\left(\theta_{i}\right)\left[\cos(\theta_{i}-\theta_{j})+\cos(\varphi_{i}-\theta_{j})\right]+J_{ji}\left(\theta_{j}\right)\left[\cos(\theta_{j}-\theta_{i})+\cos(\varphi_{j}-\theta_{i})\right]\right\}.$
Thus, we recover the Hamiltonian introduced in the Main Text, cf. Eq. (2). Based on the above result, we can derive the dynamics of $\theta_{i}$ and $\varphi_{i}$
$\dot{\theta}_{i}=\left\{\theta_{i},\mathcal{H}\right\}$ $=$ $\sum_{j\in\langle ij\rangle}J_{ij}\left(\theta_{i}\right)\sin(\varphi_{i}-\theta_{j}),$
$\dot{\varphi}_{i}=\left\{\varphi_{i},\mathcal{H}\right\}$ $=$ $\sum_{j\in\langle ij\rangle}\big{\{}-J_{ij}(\theta_{i})\sin(\theta_{i}-\theta_{j})$ (54)
$+\partial_{\theta_{i}}J_{ij}(\theta_{i})\left[\cos(\theta_{i}-\theta_{j})+\cos(\varphi_{i}-\theta_{j})\right]-J_{ji}(\theta_{j})\left[\sin(\theta_{i}-\theta_{j})+\sin(\theta_{i}-\varphi_{j})\right]\big{\}}.$
Now, let us define the constraint $G_{i}$ in phase space as the difference between the degrees of freedom $\theta_{i}$ and $\varphi_{i}$,
$G_{i}=\theta_{i}-\varphi_{i}.$ (55)
Using Eq. (B.1), the equation of motion of the constraint is given by
$\frac{\mathrm{d}G_{i}}{\mathrm{d}t}$ $=$ $\{G_{i},\mathcal{H}\}=\sum_{j}\partial_{\theta_{j}}G_{i}\partial_{\varphi_{j}}\mathcal{H}-\partial_{\varphi_{j}}G_{i}\partial_{\theta_{j}}\mathcal{H}=\dot{\theta}_{i}-\dot{\varphi}_{i}$ (56)
$=$ $\sum_{j\in\langle ij\rangle}J_{ij}\left(\theta_{i}\right)\sin(\varphi_{i}-\theta_{j})-\sum_{j\in\langle ij\rangle}\big{\{}J_{ij}(\theta_{i})\sin(\theta_{i}-\theta_{j})$
$+\partial_{\theta_{i}}J_{ij}(\theta_{i})\left[\cos(\theta_{i}-\theta_{j})+\cos(\varphi_{i}-\theta_{j})\right]-J_{ji}(\theta_{j})\left[\sin(\theta_{i}-\theta_{j})+\sin(\theta_{i}-\varphi_{j})\right]\big{\}}.$
When the degrees of freedom lie on the constraint surface, i.e., $G_{i}=\pi$, the following relations hold
$\cos\left(\theta_{i}-\theta_{j}\right)+\cos\left(\varphi_{i}-\theta_{j}\right)$ $=$ $0,$
$\sin\left(\theta_{i}-\theta_{j}\right)+\sin\left(\theta_{i}-\varphi_{j}\right)$ $=$ $0$ (57)
This results in the value of Eq. (56) being zero. Therefore, the constraint is preserved by the dynamics. Consequently, similar to the discussion in Eq. (19), if initially the system is under the constraint, $G_{i}(t=0)=\pi$, the constraint is preserved at all times: $G_{i}(t)=\pi$.
Under the constraint, the equations of motion are given by
$\dot{\theta}_{i}$ $=$ $\sum_{j\in\langle ij\rangle}J_{ij}\left(\theta_{i}\right)\sin(\varphi_{i}-\theta_{j})=-\sum_{j\in\langle ij\rangle}J_{ij}\left(\theta_{i}\right)\sin(\theta_{i}-\theta_{j}),$
$\dot{\varphi}_{i}$ $=$ $\sum_{j\in\langle ij\rangle}-J_{ij}(\theta_{i})\sin(\theta_{i}-\theta_{j})=-\sum_{j\in\langle ij\rangle}J_{ij}\left(\theta_{i}\right)\sin(\varphi_{i}-\varphi_{j}).$ (58)
Thus, the application of the constraint to the Hamiltonian embedding allows us to retrieve the overdamped nonreciprocal dynamics for the original degrees of freedom $\theta_{i}$.
B.2 Undamped nonreciprocal systems
Inspired by the approach in the previous subsection, we introduce a general procedure to construct the Hamiltonian embedding for undamped systems with nonreciprocal interactions. We will provide the explicit form of the Hamiltonian for the vision-cone XY spins, and explain how the dynamics naturally emerge from the constraint.
We first introduce the Lagrangian. It is composed of two terms: the first one is the quadratic kinetic energy $\mathcal{K}=\sum_{i}(\frac{1}{2}\dot{\theta}_{i}^{2}-\frac{1}{2}\dot{\varphi}_{i}^{2})$ and the second one is the interaction energy $\mathcal{U}$ (see below for an explanation of the negative kinetic energy for $\varphi$). The explicit form of the Lagrangian for undamped nonreciprocal systems is given by
$\mathcal{L}$ $=$ $\mathcal{K}-\mathcal{U}$ (59)
$=$ $\sum_{i}\left(\frac{1}{2}\dot{\theta}_{i}^{2}-\frac{1}{2}\dot{\varphi}_{i}^{2}\right)+\sum_{\langle ij\rangle}\left\{J_{ij}\left(\theta_{i}\right)\left[\cos(\theta_{i}-\theta_{j})+\cos(\varphi_{i}-\theta_{j})\right]+J_{ji}\left(\theta_{j}\right)\left[\cos(\theta_{j}-\theta_{i})+\cos(\varphi_{j}-\theta_{i})\right]\right\}.$
From this Lagrangian, we can derive the conjugate degree of freedom (i.e., the canonical momenta) associated with $\theta_{i}$ and $\varphi_{i}$:
$L_{i}^{\theta}=\frac{\partial\mathcal{L}}{\partial\dot{\theta}_{i}}=\dot{\theta}_{i},\qquad L_{i}^{\varphi}=\frac{\partial\mathcal{L}}{\partial\dot{\varphi}_{i}}=-\dot{\varphi}_{i},$ (60)
which leads to the Poisson bracket relations $\{\theta_{i},L_{i}^{\theta}\}=\delta_{ij}$ and $\{\varphi_{i},L_{i}^{\varphi}\}=\delta_{ij}$. Unlike the overdamped case, here we end up with two spins; for each, the phase space is a semi-compact cylinder, since the original coordinate angle is compact but angular momentum is not.
Using the canonical momenta, we can construct the corresponding Hamiltonian $\mathcal{H}$ from the Lagrangian $\mathcal{L}$ using a Legendre transform:
$\mathcal{H}$ $=$ $\sum_{i}\left(L_{i}^{\theta}\dot{\theta}_{i}+L_{i}^{\varphi}\dot{\varphi}_{i}\right)-\mathcal{L}=\mathcal{K}+\mathcal{U}$
$=$ $\frac{1}{2}\sum_{i}\left[\left({L_{i}^{\theta}}\right)^{2}-\left({L_{i}^{\varphi}}\right)^{2}\right]-\sum_{\langle ij\rangle}\left\{J_{ij}\left(\theta_{i}\right)\left[\cos(\theta_{i}-\theta_{j})+\cos(\varphi_{i}-\theta_{j})\right]+J_{ji}\left(\theta_{j}\right)\left[\cos(\theta_{j}-\theta_{i})+\cos(\varphi_{j}-\theta_{i})\right]\right\}.$
An unusual feature of the kinetic energy $\mathcal{K}$ is that it is negative-definite for the auxiliary degree of freedom $\varphi_{j}$. As elaborated on above, mathematically this is required to maintain the constraint $\theta_{i}(t)=\varphi_{i}(t)+\pi$ at all times.
Based on the above result, we derive the dynamics of $\theta_{i}$, $\varphi_{i}$, $L_{i}^{\theta}$, and $L_{i}^{\varphi}$ as follows:
$\dot{\theta}_{i}=\left\{\theta_{i},\mathcal{H}\right\}$ $=$ ${L_{i}^{\theta}},$
$\dot{\varphi}_{i}=\left\{\varphi_{i},\mathcal{H}\right\}$ $=$ $-{L_{i}^{\varphi}},$
${\dot{L}_{i}^{\theta}}=\left\{{L_{i}^{\theta}},\mathcal{H}\right\}$ $=$ $-\sum_{j\in\langle ij\rangle}\big{\{}J_{ij}(\theta_{i})\sin(\theta_{i}-\theta_{j})$
$+\partial_{\theta_{i}}J_{ij}(\theta_{i})\left[\cos(\theta_{i}-\varphi_{j})+\cos(\theta_{i}-\theta_{j})\right]+J_{ji}(\theta_{j})\left[\sin(\theta_{i}-\theta_{j})+\sin(\theta_{i}-\varphi_{j})\right]\big{\}},$
${\dot{L}_{i}^{\varphi}}=\left\{{L_{i}^{\varphi}},\mathcal{H}\right\}$ $=$ $-\sum_{j\in\langle ij\rangle}J_{ij}(\theta_{i})\sin(\varphi_{i}-\theta_{j}).$ (62)
Now let us introduce two sets of constraints $G^{(1)}_{i}$ and $G^{(2)}_{i}$. The first set represents the difference between the degrees of freedom $\theta_{i}$ and $\varphi_{i}$,
$G^{(1)}_{i}=\theta_{i}-\varphi_{i},$ (63)
and the second concerns the sum of angular momenta
${G^{(2)}_{i}=L_{i}^{\theta}+L_{i}^{\varphi}.}$ (64)
Using Eq. (B.2), the equations of motion of the constraints, $G^{(1)}_{i}$ and $G^{(2)}_{i}$, are given by
$\frac{\mathrm{d}G^{(1)}_{i}}{\mathrm{d}t}$ $=$ $\left\{G^{(1)}_{i},\mathcal{H}\right\}=L_{i}^{\theta}+L_{i}^{\varphi}=G^{(2)}_{i}$ (65)
$\frac{\mathrm{d}G^{(2)}_{i}}{\mathrm{d}t}$ $=$ $\left\{G^{(2)}_{i},\mathcal{H}\right\}=\dot{L}_{i}^{\theta}+\dot{L}_{i}^{\varphi}$ (66)
$=$ $-\sum_{j\in\langle ij\rangle}J_{ij}(\theta_{i})\sin(\theta_{i}-\theta_{j})-\sum_{j\in\langle ij\rangle}\big{\{}J_{ij}(\theta_{i})\sin(\varphi_{i}-\theta_{j})$
$+\partial_{\theta_{i}}J_{ij}(\theta_{i})\left[\cos(\theta_{i}-\varphi_{j})+\cos(\theta_{i}-\theta_{j})\right]+J_{ji}(\theta_{j})\left[\sin(\theta_{i}-\theta_{j})+\sin(\theta_{i}-\varphi_{j})\right]\big{\}}.$
When the system is on the constraint surface in phase space, defined by $G^{(1)}_{i}=\pi$ and $G^{(2)}_{i}=0$, the time derivatives of the constraints are all zero. In particular, the constraint $G^{(2)}_{i}=0$, which represents the opposite direction of the angular momenta, will make Eq. (65) evaluate to zero:
$\frac{\mathrm{d}G^{(1)}_{i}}{\mathrm{d}t}=0.$ (67)
Simultaneously, the constraint $G^{(1)}_{i}(0)=\pi$, which represents the opposite direction of the spins, leads to the following condition:
$\cos\left(\theta_{i}-\theta_{j}\right)+\cos\left(\varphi_{i}-\theta_{j}\right)$ $=$ $0,$
$\sin\left(\theta_{i}-\theta_{j}\right)+\sin\left(\varphi_{i}-\theta_{j}\right)$ $=$ $0$ (68)
This results in the value of Eq. (66) being zero:
$\frac{\mathrm{d}G^{(2)}_{i}}{\mathrm{d}t}=0.$ (69)
Similar to the overdamped case above, these constraints hold at all times.
Under the constraints, the equations of motion are given by
$\dot{\theta}_{i}=\left\{\theta_{i},\mathcal{H}\right\}$ $=$ ${L_{i}^{\theta}},$
$\dot{\varphi}_{i}=\left\{\varphi_{i},\mathcal{H}\right\}$ $=$ $-{L_{i}^{\varphi}},$
${\dot{L}_{i}^{\theta}}=\left\{{L_{i}^{\theta}},\mathcal{H}\right\}$ $=$ $-\sum_{j\in\langle ij\rangle}J_{ij}(\theta_{i})\sin(\theta_{i}-\theta_{j})$
${\dot{L}_{i}^{\varphi}}=\left\{{L_{i}^{\varphi}},\mathcal{H}\right\}$ $=$ $\sum_{j\in\langle ij\rangle}J_{ij}(\theta_{i})\sin(\varphi_{i}-\varphi_{j}).$ (70)
We can also write them in the form of the second-order time derivative of $\theta_{i}$ and $\varphi_{i}$,
$\ddot{\theta}_{i}$ $=$ $-\sum_{j\in\langle ij\rangle}J_{ij}(\theta_{i})\sin(\theta_{i}-\theta_{j}),$
$\ddot{\varphi}_{i}$ $=$ $-\sum_{j\in\langle ij\rangle}J_{ij}(\theta_{i})\sin(\varphi_{i}-\varphi_{j}).$ (71)
Thus, the application of the constraints to the Hamiltonian embedding allows us to retrieve the undamped nonreciprocal dynamics for the original degree of freedom $\theta_{i}$.
The occurrence of negative kinetic energy for the auxiliary degrees of freedom can be explained as follows: Consider first the free system where the degrees of freedom are noninteracting, $\mathcal{U}=0$, and both terms in the kinetic energy are positive; then, by conservation of angular momentum $L^{\varphi}_{j}+L^{\theta}_{j}=0$, the $\theta$ and $\varphi$ degrees of freedom will rotate opposite to one another, thus violating the constraint that pins them to move together (i.e., in the same direction). To enforce the latter without over-constraining the dynamics of the resulting joint motion, a negative kinetic energy is required. Returning to the Hamiltonian (B.2), more formally, as discussed above, we need to construct a constraint as $\theta_{i}(t)=\varphi_{i}(t)+\pi$. Taking two time derivatives, we have $\ddot{\theta}_{i}(t)=\ddot{\varphi}_{i}(t)$. If the interactions between the original and the auxiliary degrees of freedom are defined as in Eq. (21), the torques experienced by the original spin and the auxiliary spin will be oppositely directed under the imposed constraint, as given by $-\sum_{j\in\langle ij\rangle}J_{ij}\sin(\theta_{i}-\theta_{j})=-\big{[}-\sum_{j\in\langle ij\rangle}J_{ij}\sin(\varphi_{i}-\theta_{j})\big{]}$. To maintain the opposite sign in the EOM of the angular momenta, the kinetic energy for the auxiliary degree of freedom has a negative sign, $-\frac{1}{2}\dot{\varphi}_{i}^{2}$. Such negative kinetic energy terms for the auxiliary degree of freedom have been discussed in Ref. [56] where they arise as a result of reformulating the action principle as an initial- (rather than a boundary-) value problem. Similarly, in the programmed robotic metamaterials example from Sec. A.4, we used the kinetic term $\frac{1}{2}\sum_{i}\left(p_{i}^{u}\right)^{2}-\left(p_{i}^{w}\right)^{2}$ to construct the Hamiltonian. Such anisotropic kinetic energy terms have also been encountered before in the description of azimuthal curl forces [76].
As we showed at the beginning of Appendix A, one can extend the arguments of this section to more general nonreciprocal systems whose equations of motion feature both inertial and dissipative terms.
Appendix C Equivalence between nonreciprocal Langevin dynamics and Glauber Monte Carlo dynamics
This section demonstrates that the steady-state density distributions resulting from the original nonreciprocal Langevin dynamics and from the constrained Glauber dynamics based on the Hamiltonian that we propose are identical. In Sec. C.1, we first derive a Fokker-Planck-like equation for the phase-space density distribution obtained from Langevin dynamics; then, in Sec. C.2, we prove that it is equivalent to the corresponding distribution obtained using constrained Glauber dynamics.
Below, we consider for concreteness the nonreciprocal Hamiltonian for the vision-cone interactions (see Main Text for the details)
$H=H_{SS}+H_{Sa}=-\sum_{\langle ij\rangle}\left\{\left[J_{ij}(\theta_{i})+J_{ji}(\theta_{j})\right]\cos(\theta_{i}-\theta_{j})+J_{ij}(\theta_{i})\cos(\theta_{j}-\varphi_{i})+J_{ji}(\theta_{j})\cos(\theta_{i}-\varphi_{j})\right\},$ (72)
although the procedure is general and applies to other models (see Sec. A).
C.1 Langevin steady-state distribution
Recall that imposing the constraint $\theta_{i}(t)-\varphi_{i}(t)=\pi$, and adding the same noise term for both the original and the auxiliary degrees of freedom results in the overdamped Langevin equations
$\dot{\theta_{i}}$ $=$ $\left\{\theta_{i},H\right\}+\sqrt{2T}\eta_{i}(t)=\partial_{\varphi_{i}}H|_{\text{constr.}}+\sqrt{2T}\eta_{i}(t),$
$\dot{\varphi}_{i}$ $=$ $\left\{\varphi_{i},H\right\}+\sqrt{2T}\eta_{i}(t)=-\partial_{\theta_{i}}H|_{\text{constr.}}+\sqrt{2T}\eta_{i}(t),$ (73)
with conjugate degrees of freedom $\{\theta_{i},\varphi_{j}\}{=}\delta_{ij}$. To maintain consistency with the traditional form of Langevin dynamics and represent the equations of motion on the constraint surface, the notation $\partial_{\theta_{i}}H|_{\text{constr.}}$ is introduced to replace the Poisson bracket. Here $\eta_{i}(t)$ denotes white noise with $\langle\eta_{i}(t)\rangle=0$ and $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$, with $T$ the temperature of the bath, where the average is taken over the phase-space density $\rho(\theta_{i}(t),\varphi_{i}(t))$. Note that the constraint implies that each pair $(\theta_{i},\varphi_{i})$ experiences the same noise realization.
Below, we first consider steady-state distributions (i.e., no explicit $t$-dependence in $\rho$, Sec. C.1.1), and then generalize the result to non-stationary distributions in Sec. C.1.2.
C.1.1 Steady-state distribution
To obtain the steady-state phase-space density distribution $\rho(\theta_{i}(t),\varphi_{i}(t))$ under the above dynamics, we follow the standard procedure [54, 55] and introduce a test function $f(\theta_{i}(t),\varphi_{i}(t))$ over phase space. The change of the test function within a small time window $\Delta t$ is to leading order
$f\left(\theta_{i}(t+\Delta t),{\varphi}_{i}(t+\Delta t)\right)-f\left(\theta_{i}(t),{\varphi}_{i}(t)\right)$ (74)
$=$ $\sum_{i}\frac{\partial f}{\partial\theta_{i}}\dot{\theta_{i}}\Delta t+\frac{\partial f}{\partial\varphi_{i}}\dot{\varphi}_{i}\Delta t+\frac{\partial^{2}f}{2\partial\theta_{i}\partial\theta_{i}}(\dot{\theta_{i}}\Delta t)^{2}+\frac{\partial^{2}f}{2\partial\varphi_{i}\partial\varphi_{i}}(\dot{\varphi}_{i}\Delta t)^{2}+\frac{\partial^{2}f}{\partial\varphi_{i}\partial\theta_{i}}\dot{\varphi_{i}}\dot{\theta_{i}}(\Delta t)^{2}$
$+\sum_{i\neq j}\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\dot{\theta_{i}}\dot{\theta_{j}}(\Delta t)^{2}+\frac{\partial^{2}f}{\partial\varphi_{i}\partial\varphi_{j}}\dot{\varphi_{i}}\dot{\varphi_{j}}(\Delta t)^{2}+\cdots\ .$
For steady-state distributions, at long times the expectation value of the test function should be time-independent:
$\langle f\left(t+\Delta t\right)\rangle=\langle f\left(t\right)\rangle,$ (75)
which can be written infinitesimally as
$\int\frac{\mathrm{d}f\left(\theta_{i},\varphi_{i}\right)}{\mathrm{d}t}\rho(\theta_{i},\varphi_{i})\prod_{j}\mathrm{d}{\theta_{j}}\mathrm{d}{\varphi_{j}}=\lim_{\Delta t\to 0}\int\frac{f\left(\theta_{i}(t+\Delta t),{\varphi}_{i}(t+\Delta t)\right)-f\left(\theta_{i}(t),{\varphi}_{i}(t)\right)}{\Delta t}\rho(\theta_{i},\varphi_{i})\prod_{j}\mathrm{d}{\theta_{j}}\mathrm{d}{\varphi_{j}}=0.$ (76)
To proceed with the derivation, we write the noise terms in Eq. (C.1) as a Wiener process $W_{i}(t)$:
$\dot{\theta}_{i}\Delta t=\partial_{\varphi_{i}}H|_{\text{constr.}}\Delta t+\sqrt{2T}\Delta W_{i}(t),$
$\dot{\varphi}_{i}\Delta t=-\partial_{\theta_{i}}H|_{\text{constr.}}\Delta t+\sqrt{2T}\Delta W_{i}(t),$ (77)
where $\langle\Delta W_{i}(t)\rangle=0$ and $\langle\Delta W_{i}(t)\Delta W_{j}(t^{\prime})\rangle=\Delta t\delta(t-t^{\prime})\delta_{ij}$.
In the following, we substitute Eq. (74) in Eq. (76). Focusing on the $\theta_{i}$ terms, the first-order contribution can be evaluated as
$\int\frac{\partial f}{\partial\theta_{i}}\dot{\theta_{i}}\Delta t\;\rho\left(\theta_{i},\varphi_{i}\right)\mathrm{d}\mathbf{\theta}_{i}\mathrm{d}\mathbf{\varphi}_{i}$ $=$ $\int\frac{\partial f}{\partial\theta_{i}}\left[\frac{\partial H}{\partial\varphi_{i}}\Bigg{|}_{\text{constr.}}\Delta t+\sqrt{2T}\Delta W_{i}(t)\right]\rho\left(\theta_{i},\varphi_{i}\right)\mathrm{d}{\theta_{i}}\mathrm{d}{\varphi_{i}}$ (78)
$=$ $\Delta t\int\frac{\partial f}{\partial\theta_{i}}\frac{\partial H}{\partial\varphi_{i}}\Bigg{|}_{\text{constr.}}\rho\left(\theta_{i},\varphi_{i}\right)\mathrm{d}{\theta_{i}}\mathrm{d}{\varphi_{i}},$
where we used the equation of motion and the property $\langle\Delta W_{i}(t)\rangle=0$ in the last line. Similarly, the second order in $\Delta t$ gives
$\int\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\dot{\theta_{i}}\dot{\theta_{j}}\Delta t^{2}\rho\left(\theta_{i},\varphi_{i}\right)\mathrm{d}{\theta_{i}}\mathrm{d}{\varphi_{i}}=$ (79)
$=$ $\int\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\left[\frac{\partial H}{\partial\varphi_{i}}\Bigg{|}_{\text{constr.}}\Delta t+\sqrt{2T}\Delta W_{i}(t)\right]\left[\frac{\partial H}{\partial\varphi_{j}}\Bigg{|}_{\text{constr.}}\Delta t+\sqrt{2T}\Delta W_{j}(t)\right]\rho\left(\theta_{i},\varphi_{i}\right)\mathrm{d}{\theta_{i}}\mathrm{d}{\varphi_{i}}$
$=$ $\int 2T\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\Delta W_{i}(t)\Delta W_{j}(t)\rho\left(\theta_{i},\varphi_{i}\right)\mathrm{d}\theta_{i}\mathrm{d}\varphi_{i}+\sqrt{2T}\int\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\frac{\partial H}{\partial\varphi_{j}}\Bigg{|}_{\text{constr.}}\Delta W_{i}(t)\rho\left(\theta_{i},\varphi_{i}\right)\Delta t\mathrm{d}{\theta_{i}}\mathrm{d}{\varphi_{i}}$
$+\sqrt{2T}\int\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\frac{\partial H}{\partial\varphi_{i}}\Bigg{|}_{\text{constr.}}\Delta W_{j}(t)\rho\left(\theta_{i},\varphi_{i}\right)\Delta t\mathrm{d}{\theta_{i}}\mathrm{d}{\varphi_{i}}+\Delta t^{2}\int\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\left(\frac{\partial H}{\partial\varphi_{i}}\Bigg{|}_{\text{constr.}}\frac{\partial H}{\partial\varphi_{j}}\Bigg{|}_{\text{constr.}}\right)\rho\left(\theta_{i},\varphi_{i}\right)\mathrm{d}{\theta_{i}}\mathrm{d}{\varphi_{i}}$
$=$ $\Delta t\delta_{ij}2T\int\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{j}}\rho\left(\theta_{i},\varphi_{i},t\right)\mathrm{d}\theta_{i}\mathrm{d}\varphi_{i}+O\left[\left(\Delta t\right)^{2}\right],$
where we neglect terms of order $\Delta t^{2}$ since they vanish in the infinitesimal limit $\Delta t\to 0$. The expressions for the other terms in Eq. (74) can be derived using the same reasoning.
Collecting all $\Delta t$ contributions, and plugging them back to Eq. (76) leads to the following condition
$\sum_{i}\int\prod_{j\neq i}\mathrm{d}\theta_{j}\mathrm{d}\varphi_{j}\int\mathrm{d}\theta_{i}\mathrm{d}\varphi_{i}\Bigg{[}\frac{\partial f}{\partial\theta_{i}}\left(\frac{\partial H}{\partial\varphi_{i}}\Bigg{|}_{\text{constr.}}\right)-\frac{\partial f}{\partial\varphi_{i}}\left(\frac{\partial H}{\partial\theta_{i}}\Bigg{|}_{\text{constr.}}\right)$
$+T\left(\frac{\partial^{2}f}{\partial\theta_{i}\partial\theta_{i}}+\frac{\partial^{2}f}{\partial\varphi_{i}\partial\varphi_{i}}+2\frac{\partial^{2}f}{\partial\varphi_{i}\partial\theta_{i}}\right)\Bigg{]}\rho\left(\theta_{i},\varphi_{i}\right)$ $=$ $0\,.$ (80)
Noticing that the term proportional to $T$ is a perfect square, we have
$\sum_{i}\int\prod_{j\neq i}\mathrm{d}\theta_{j}\mathrm{d}\varphi_{j}\int\mathrm{d}\theta_{i}\mathrm{d}\varphi_{i}\left[\frac{\partial f}{\partial\theta_{i}}\left(\frac{\partial H}{\partial\varphi_{i}}\Bigg{|}_{\text{constr.}}\right)-\frac{\partial f}{\partial\varphi_{i}}\left(\frac{\partial H}{\partial\theta_{i}}\Bigg{|}_{\text{constr.}}\right)+T\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)^{2}f\right]\rho\left(\theta_{i},\varphi_{i}\right)=0.$ (81)
Under the constraint $\theta_{i}-\varphi_{i}=\pi$, the dynamics of the $\varphi_{i}$ spins mirrors that of the $\theta_{i}$ spins. Hence, we have $\partial_{\varphi_{i}}H|_{\text{constr.}}=-\partial_{\theta_{i}}H|_{\text{constr.}}$. Plugging this into Eq. (81) leads to
$\sum_{i}\int\prod_{j\neq i}\mathrm{d}\theta_{j}\mathrm{d}\varphi_{j}\int\mathrm{d}\theta_{i}\mathrm{d}\varphi_{i}\left[-\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)f\frac{\partial H}{\partial\theta_{i}}\Bigg{|}_{\text{constr.}}+T\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)^{2}f\right]\rho\left(\theta_{i},\varphi_{i}\right)=0.$ (82)
If we integrate the $T$ term by parts and, recalling there is no boundary term for a compact phase space, we obtain
$\sum_{i}\int\prod_{j\neq i}\mathrm{d}\theta_{j}\mathrm{d}\varphi_{j}\int\mathrm{d}\theta_{i}\mathrm{d}\varphi_{i}f\left[\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)\left(\frac{\partial H}{\partial\theta_{i}}\Bigg{|}_{\text{constr.}}\rho\left(\theta_{i},\varphi_{i}\right)\right)+T\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)^{2}\rho\left(\theta_{i},\varphi_{i}\right)\right]=0.$ (83)
Since this equation for the steady state is valid for all test functions $f$, we find that the phase-space density in the steady state satisfies
$\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)\left[\frac{\partial H}{\partial\theta_{i}}\Bigg{|}_{\text{constr.}}\ \rho\left(\theta_{i},\varphi_{i}\right)+T\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)\rho\left(\theta_{i},\varphi_{i}\right)\right]=0.$ (84)
So far the derivation was generic since we did not make use of the precise form of the equations of motion. We will show in Sec. C.2 that Eq. (84) governs also the constrained Glauber dynamics.
C.1.2 Time-dependent non-stationary distribution
Before we do that, let us complete the analysis and examine the case where the density distribution exhibits explicit time dependence, $\rho(\theta_{i}(t),\varphi_{i}(t),t)$. Such situations can occur away from equilibrium where oscillating states arise.
First, we allow the test function to include explicit time dependence. Building on the same analytical framework as above, the evolution of the expectation value of the test function takes the form
$\frac{\mathrm{d}\left\langle f\right\rangle}{\mathrm{d}t}=\left\langle\frac{\partial f}{\partial t}\right\rangle+\sum_{i}\left(\left\langle\frac{\partial f}{\partial\theta_{i}}\frac{\partial H}{\partial\varphi_{i}}\Bigg{|}_{\text{constr.}}\right\rangle-\left\langle\frac{\partial f}{\partial\varphi_{i}}\frac{\partial H}{\partial\theta_{i}}\Bigg{|}_{\text{constr.}}\right\rangle+T\left\langle\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)^{2}f\right\rangle\right),$ (85)
which can be written as
$\frac{\mathrm{d}}{\mathrm{d}t}\int\prod_{j}\mathrm{d}\theta_{j}\mathrm{d}\varphi_{j}f\rho=$
$=$ $\int\prod_{j}\mathrm{d}\theta_{j}\mathrm{d}\varphi_{j}\frac{\partial f}{\partial t}\rho+\sum_{i}\int\prod_{j\neq i}\mathrm{d}\theta_{j}\mathrm{d}\varphi_{j}\int\mathrm{d}\theta_{i}\mathrm{d}\varphi_{i}\Bigg{[}\frac{\partial f}{\partial\theta_{i}}\frac{\partial H}{\partial\varphi_{i}}\Bigg{|}_{\text{constr.}}\rho-\frac{\partial f}{\partial\varphi_{i}}\frac{\partial H}{\partial\theta_{i}}\Bigg{|}_{\text{constr.}}\rho+T\int\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)^{2}f\rho\Bigg{]}.$
On the other hand, directly from the left-hand side of the equation above, we have
$\frac{\mathrm{d}}{\mathrm{d}t}\int\prod_{j}\mathrm{d}\theta_{j}\mathrm{d}\varphi_{j}f\rho=\int\prod_{j}\mathrm{d}\theta_{j}\mathrm{d}\varphi_{j}\left(\frac{\partial f}{\partial t}\rho+f\frac{\partial\rho}{\partial t}\right).$ (87)
Setting Eq.(C.1.2) equal to Eq.(87) yields
$\int\prod_{j}\mathrm{d}\theta_{j}\mathrm{d}\varphi_{j}f\frac{\partial\rho}{\partial t}=\sum_{i}\int\prod_{j\neq i}\mathrm{d}\theta_{j}\mathrm{d}\varphi_{j}\int\mathrm{d}\theta_{i}\mathrm{d}\varphi_{i}\Bigg{[}\frac{\partial f}{\partial t}\rho+\frac{\partial f}{\partial\theta_{i}}\frac{\partial H}{\partial\varphi_{i}}\Bigg{|}_{\text{constr.}}\rho-\frac{\partial f}{\partial\varphi_{i}}\frac{\partial H}{\partial\theta_{i}}\Bigg{|}_{\text{constr.}}\rho+T\int\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)^{2}f\rho\Bigg{]}\,.$
After integrating the right-hand side of this equation by parts and plugging $\partial_{\varphi_{i}}H|_{\text{constr.}}=-\partial_{\theta_{i}}H|_{\text{constr.}}$ into the above equation, we have
$\int\prod_{j}\mathrm{d}\theta_{j}\mathrm{d}\varphi_{j}f\frac{\partial\rho}{\partial t}=\int\prod_{j}\mathrm{d}\theta_{j}\mathrm{d}\varphi_{j}f\sum_{i}\left[\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)\left(\frac{\partial H}{\partial\theta_{i}}\Bigg{|}_{\text{constr.}}\rho\right)+T\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)^{2}\rho\right].$ (88)
Since this result has to hold for all test functions $f\left(\theta_{i}(t),\varphi_{i}(t),t\right)$, we find
$\frac{\partial\rho}{\partial t}=\sum_{i}\left[\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)\left(\frac{\partial H}{\partial\theta_{i}}\Bigg{|}_{\text{constr.}}\rho\right)+T\left(\frac{\partial}{\partial\theta_{i}}+\frac{\partial}{\partial\varphi_{i}}\right)^{2}\rho\right].$ (89)
This Fokker-Planck equation governs the dynamics of the phase-space distribution. As expected, for a stationary distribution, $\partial_{t}\rho=0$, it reduces to Eq. (84).
C.2 Glauber dynamics description
We now present two equivalent Glauber algorithms that sample the steady state of the Langevin dynamics. In particular, Sec. C.2.1 presents the algorithm used in the main text, while Sec. C.2.2 presents an alternative, yet equivalent, Glauber algorithm based on global balance.
C.2.1 Individual-update Glauber scheme: Detailed balance followed by imposing the constraint
In this subsection, we introduce our constrained Glauber dynamics, which is based on the Hamiltonian embedding that we propose for nonreciprocal systems. It is a Markov process with the constraint that every state $(\theta_{i},\varphi_{i})$ satisfies the condition $\theta_{i}=\varphi_{i}+\pi$. Similar to the Langevin description in the previous section, we assume that the dynamics eventually reaches a steady state, with the density distribution given by $\rho(\theta,\varphi)$.
Note that, evaluating the Hamiltonian on the constraint, we find $H(\theta_{i},\varphi_{i})=0$, which makes it difficult to define a proper Glauber update rule. Therefore, the algorithm starts by first proposing an update of only the original spin, changing its angle as $\theta\rightarrow\theta^{\prime}=\theta+\delta$. The transition rates $w(\theta\rightarrow\theta^{\prime})$ of this proposed update are calculated based on the energy difference between the original and updated state, obtained from the Hamiltonian. Subsequently, the auxiliary spin is rotated accordingly as $\varphi\rightarrow\varphi+\delta$ to enforce the mirror constraint. In practice, enforcing the constraint only in the second step allows us to define transition rates using the energy fluctuations caused by a fluctuation in $\theta$ at fixed $\varphi$. Below we show that keeping $\delta$ small ensures the agreement of the Glauber and Langevin steady state.
In the steady state, the phase-space distribution is stationary. Hence, the probability fluxes into and out of a given state balance out:
$\left[\rho(\theta_{i}+\delta,\varphi_{i}+\delta)w(\theta_{i}+\delta\rightarrow\theta_{i})-\rho(\theta_{i},\varphi_{i})w(\theta_{i}\rightarrow\theta_{i}+\delta)\right]\Big{|}_{\text{constr.}}=0\ ,$ (90)
where the specific form of the Glauber transition rate at constant $\varphi_{i}$ is
$w\left(\theta_{i}\rightarrow\theta_{i}^{\prime}\right)=\frac{1}{2}\left(1-\tanh\frac{H(\theta^{\prime}_{i},\varphi_{i})-H(\theta_{i},\varphi_{i})}{2T}\right)\,.$ (91)
Note the explicit presence of the constraint in Eq. (90), corresponding to Step 2 from the main text. We emphasize that this ad-hoc procedure violates detailed balance – a prerequisite for having a nonequilibrium steady state.
Next, we approximate Eq. (90) for small $\delta$ as
$\Bigg{\{}\left[\rho(\theta_{i},\varphi_{i})+\delta\times\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\rho(\theta_{i},\varphi_{i})\right]\frac{1}{2}\left(1-\tanh\frac{-\delta\times\partial_{\theta_{i}}H}{2T}\right)-\rho(\theta_{i},\varphi_{i})\frac{1}{2}\left(1-\tanh\frac{+\delta\times\partial_{\theta_{i}}H}{2T}\right)\Bigg{\}}\Bigg{|}_{\text{constr.}}=0.$
Performing a Taylor expansion on the above equation keeping terms to first order in $\delta$ and multiplying by $T$ yields
$\delta\Big{\{}T\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\rho(\theta_{i},\varphi_{i})+\rho(\theta_{i},\varphi_{i})\partial_{\theta_{i}}H\Big{\}}\Big{|}_{\text{constr.}}=0.$ (92)
Since this condition must be satisfied for any $\delta$, we arrive at the following equation for the phase-space distribution function $\rho$
$T\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\rho(\theta_{i},\varphi_{i})+\rho(\theta_{i},\varphi_{i})\partial_{\theta_{i}}H\Big{|}_{\text{constr.}}$ $=$ $0$ (93)
Thus, we find that Eq. (93) satisfies Eq. (84) derived above. Therefore, the long-time steady state of the constrained Glauber dynamics reproduces the steady state of the Langevin dynamics.
C.2.2 Joint-update Glauber scheme: Global balance
Let us now discuss an alternative Glauber algorithm, which leads to the same result. Consider again a Markov process where every state $(\theta_{i},\varphi_{i})$ satisfies the constraint $\theta_{i}=\varphi_{i}+\pi$. This time, the constraint is enforced by simultaneously updating the original and auxiliary spins, which change their angles by the same amount $\delta$ as $\theta\rightarrow\theta^{\prime}=\theta+\delta$ and $\varphi\rightarrow\varphi^{\prime}=\varphi+\delta$.
Assuming that the dynamics reaches a nonequilibrium steady state, it satisfies global balance [38]
$\int_{-\pi}^{\pi}\left[\rho(\theta_{i}+\delta,\varphi_{i}+\delta)W(\theta_{i}+\delta,\varphi_{i}+\delta\rightarrow\theta_{i},\varphi_{i})-\rho(\theta_{i},\varphi_{i})W(\theta_{i},\varphi_{i}\rightarrow\theta_{i}+\delta,\varphi_{i}+\delta)\right]\mathrm{d}\delta=0,$ (94)
where the constraint is built-in. Note the additional freedom we have in choosing the precise form of the rates $W$: so long as they satisfy global balance, they can be used to sample the steady-state distribution.
[…]
$w\left(\theta_{i}\rightarrow\theta_{i}^{\prime},\varphi_{i}\rightarrow\varphi_{i}^{\prime}\right)=\frac{1}{2}\left(1-\tanh\left(\frac{H\left(\theta^{\prime},\varphi\right)-H\left(\theta,\varphi^{\prime}\right)}{4T}\right)\right)\,.$ (95)
In other words, we compare the change in energy resulting from the update of each variable $\theta,\varphi$, separately. In particular, we change the angle of the original spin as $\theta\rightarrow\theta^{\prime}{=}\theta+\delta$. The resulting energy difference between the two states is given by $H(\theta^{\prime}_{i},\varphi_{i})-H(\theta_{i},\varphi_{i})$. We also compute the energy difference when the corresponding auxiliary spin is changed under the constraint, i.e., $\varphi\rightarrow\varphi^{\prime}{=}\varphi+\delta$, with the energy difference $H(\theta_{i},\varphi^{\prime}_{i})-H(\theta_{i},\varphi_{i})$. This gives a total change of $H\left(\theta^{\prime},\varphi\right)-H\left(\theta,\varphi^{\prime}\right)$. There is an additional factor of $1/2$ in the argument of the $\tanh$-function, compared to the previous section’s rates, which takes into account the update of both degrees of freedom.
For small $\delta\in[-\Delta,\Delta]$ we can approximate Eq. (94) as
$\int_{-\Delta}^{\Delta}\Bigg{\{}\left[\rho(\theta_{i},\varphi_{i})+\delta\times\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\rho(\theta_{i},\varphi_{i})+\frac{1}{2}\delta^{2}\times\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)^{2}\rho(\theta_{i},\varphi_{i})\right]$
×$\times\frac{1}{2}\left(1-\tanh\frac{-2\left(\partial_{\theta_{i}}H-\partial_{\varphi_{i}}H\right)\times\delta-\left(\partial_{\theta_{i}}^{2}H-\partial_{\varphi_{i}}^{2}H\right)\times\delta^{2}}{8T}\right)$
$-\rho(\theta_{i},\varphi_{i})\frac{1}{2}\left(1+\tanh\frac{-2\left(\partial_{\theta_{i}}H-\partial_{\varphi_{i}}H\right)\times\delta-\left(\partial_{\theta_{i}}^{2}H-\partial_{\varphi_{i}}^{2}H\right)\times\delta^{2}}{8T}\right)\Bigg{\}}\Bigg{|}_{\text{constr.}}\ \mathrm{d}\delta=0.$ (96)
Next, we perform a Taylor expansion on the above equation. The first-order contribution vanishes immediately due to the symmetry of the integrand:
$\frac{1}{2}\Big{\{}T\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\rho(\theta_{i},\varphi_{i})+\frac{1}{2}\rho(\theta_{i},\varphi_{i})\left(\partial_{\theta_{i}}H-\partial_{\varphi_{i}}H\right)\Big{\}}\Big{|}_{\text{constr.}}\;\int_{-\Delta}^{\Delta}\delta\;\mathrm{d}\delta=0.$ (97)
The second-order contribution in $\delta$ does not vanish. Multiplying by $T$, we have
$\frac{1}{4}\Big{\{}\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\left[T\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\rho(\theta_{i},\varphi_{i})+\frac{1}{2}\rho(\theta_{i},\varphi_{i})\left(\partial_{\theta_{i}}H-\partial_{\varphi_{i}}H\right)\right]\Big{\}}\Big{|}_{\text{constr.}}\ \int_{-\Delta}^{\Delta}\delta^{2}\mathrm{d}\delta$ $=$ $0$ (99)
Rearranging and cancelling the constants, we arrive at the following equation for the phase-space distribution function $\rho$
$\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\left[T\left(\partial_{\theta_{i}}+\partial_{\varphi_{i}}\right)\rho(\theta_{i},\varphi_{i})+\frac{1}{2}\rho(\theta_{i},\varphi_{i})\left(\partial_{\theta_{i}}H-\partial_{\varphi_{i}}H\right)\right]\Big{|}_{\text{constr.}}\ =0.$ (100)
Multiplying by a test function $f(\theta_{i},\varphi_{i})$ and integrating over phase space, we find that Eq. (100) is equivalent to Eq. (84) since, under the constraint, we have $-\partial_{\varphi_{i}}H=\partial_{\theta_{i}}H$.
Both of these derivations show that the constrained Glauber algorithms proposed here can be used to sample steady states of nonreciprocal systems efficiently. In Figure 6, we test these algorithms against the original Langevin dynamics. The average magnetization and specific heat obtained from Monte-Carlo simulations of constrained Glauber dynamics using the joint update (red) and individual update (blue) exhibit behaviors similar to those obtained from Langevin dynamics. All three methods yield the same critical temperature, $T_{c}/J=0.79\pm 0.03$ (joint update and individual update) and $T_{c}/J=0.78\pm 0.02$ (Langevin) within statistical uncertainty, for the phase transition, cf. Sec. D. Notice that the average magnetization and standard error obtained from these three methods generally agree well, except near the critical temperature, where fluctuations are stronger due to criticality.
Unless explicitly stated otherwise, we use the individual-update Glauber algorithm from Sec. C.2.1 for the numerical simulations in this work.
Figure 6: Average magnetization $\langle m\rangle$ (a) and specific heat $C_{v}$ (b) obtained from Monte-Carlo simulations of constrained Glauber dynamics using the joint update (red, Sec. C.2.2), and individual update (blue, Sec. C.2.1), as well as simulations of the original Langevin dynamics (green) as functions of temperature $T$. All methods yield the same critical temperature within statistical uncertainty, $T_{c}=0.79\pm 0.03J$ (joint update and individual update) and $T_{c}=0.78\pm 0.02J$ (Langevin). The remaining parameter values are the same as in Fig. 3.
Appendix D Determining the critical temperature
To determine the critical temperature within the statistical uncertainty, $T_{c}\pm\Delta T_{c}$, for a given algorithm, we first conduct $N$ independent simulations (runs) of the dynamics for fixed values $T$ of the reservoir temperature. For each run $i$, we determine the transition temperature $T_{c,i}$ from the peak of the specific heat. The mean and standard deviation over the $N$ independent runs are subsequently computed from
$T_{c}$ $=$ $\frac{1}{N}\sum_{i=1}^{N}T_{c,i},\qquad\qquad\Delta T_{c}=\frac{1}{N}\sqrt{\sum_{i=1}^{N}\left(T_{c,i}-T_{c}\right)^{2}}.$ (101)
Appendix E Supplementary results for the Monte-Carlo simulations
This section presents supplementary results for the Monte-Carlo simulations presented in the Main Text. In Sec. E.1, we detail the implementation of the Glauber algorithm under the proposed constraints, derived in Sec. C.2.1. In Sec. E.2, we present numerical results for the individual-update Glauber dynamics that we propose, for Glauber dynamics with selfish energy as proposed in previous works [30, 29, 50], and for Langevin dynamics across various system sizes. Additionally, we examine the outcomes of Glauber dynamics with the Hamiltonian dropping the auxiliary spin (Eq. (104)), and highlight the significant differences from the behavior observed in the constrained Glauber dynamics. Finally, in Sec. E.4, we explore a different vision-cone interaction strength proposed in Ref. [45] and compare the constrained Glauber dynamics with selfish energy dynamics [45] and Langevin dynamics: We find numerically identical behavior obtained from the constrained Glauber dynamics and the Langevin dynamics, whereas the selfish energy approach results in a different behavior.
E.1 Individual-update Glauber algorithm with constraint
In this subsection, we introduce the details of the constrained Glauber algorithm from Sec. C.2.1 that we use in the Main Text. The Hamiltonian is introduced in Eq. (72). We use $\Psi^{(n)}$ to represent the system configuration $\Psi^{(n)}=(\theta_{1}^{(n)},\theta_{2}^{(n)},...\theta_{i}^{(n)},...\theta_{N}^{(n)},\varphi_{1}^{(n)},\varphi_{2}^{(n)},...\varphi_{i}^{(n)},...\varphi_{N}^{(n)})$, where $n$ denotes the iteration step of the Monte-Carlo simulation. The algorithm proceeds as follows:
  1. 1.
    Initial Configuration: Set initial values for the original and auxiliary spins in a uniform configuration, with $\theta_{i}^{(0)}=0$ and $\varphi_{i}^{(0)}=\pi$. This initial configuration satisfies the constraint.
  2. 2.
    Single Spin Update: Propose a new configuration by selecting an original spin $\theta_{i}^{(0)}$ randomly and rotating it to a new angle $\theta_{i}^{(0)}\rightarrow\theta_{i}^{(0)}+\delta$, where $\delta$ is drawn from a uniform distribution between $[-\Delta,\Delta]$. As anticipated in Sec. C.2.1, we find empirically that the smaller $\Delta$, the better the agreement with the Langevin steady state.
  3. 3.
    Energy Change Calculation: The energy difference between the initial and the proposed configurations is given by $\Delta H=H(\theta_{i}^{(0)}+\delta,\varphi_{i}^{(0)})-H(\theta_{i}^{(0)},\varphi_{i}^{(0)})$. For the XY spins with vision-cone interactions, it is given by
    $\Delta H$ $=$ $-\sum_{j\in\langle ij\rangle}\bigg{\{}J_{ij}(\theta_{i}^{(0)}+\delta)\left[\cos(\theta_{i}^{(0)}+\delta-\theta_{j}^{(0)})+\cos(\varphi_{i}^{(0)}-\theta_{j}^{(0)})\right]$
    $\qquad\qquad+J_{ji}(\theta_{j}^{(0)})\left[\cos(\theta_{j}^{(0)}-\theta_{i}^{(0)}-\delta)+\cos(\varphi_{j}^{(0)}-\theta_{i}^{(0)}-\delta)\right]\bigg{\}}$
    $+\sum_{j\in\langle ij\rangle}\bigg{\{}J_{ij}(\theta_{i}^{(0)})\left[\cos(\theta_{i}^{(0)}-\theta_{j}^{(0)})+\cos(\varphi_{i}^{(0)}-\theta_{j}^{(0)})\right]+J_{ji}(\theta_{j}^{(0)})\left[\cos(\theta_{j}^{(0)}-\theta_{i}^{(0)})+\cos(\varphi_{j}^{(0)}-\theta_{i}^{(0)})\right]\bigg{\}}.$
  4. 4.
    Transition Probability: Calculate the transition probability $w$ using the individual-update Glauber transition rates at temperature $T$:
    $w=\frac{1}{2}\left(1-\tanh\frac{\Delta H}{2T}\right).$ (103)
  5. 5.
    Acceptance/Rejection Step: To accept or reject the proposed new configuration, draw a random number $r$ from a uniform distribution between 00 and $1$. If it is less than the transition probability, i.e., $r<w$, accept the new configuration by updating the selected spin as $\theta_{i}^{(1)}=\theta_{i}^{(0)}+\delta$. Otherwise, reject the update and leave the spin unchanged. Note that all other spins remain unchanged $\theta_{j}^{(1)}=\theta_{j}^{(0)}$, for all $j\neq i$.
  6. 6.
    Enforce Constraint: Update the auxiliary spins, $\varphi_{i}^{(1)}=\theta_{i}^{(1)}+\pi$ to enforce the constraint. We thus obtain the next-step configuration $\Psi^{(1)}$.
  7. 7.
    Repeat: Iterate the procedure for a large number of steps. In our numerical simulations, we perform $N^{5}$ iterations where $N$ is the system size.
We note in passing that a similar algorithm can be constructed for the joint-update scheme in a straightforward way, following the analysis in Sec. C.2.2.
Figure 7: Average magnetization $\langle m\rangle$ (a,b,c) and specific heat $C_{v}$ (d,e,f) obtained from numerical simulations of Langevin dynamics, individual-update constrained Glauber dynamics, and selfish-energy Glauber dynamics as functions of temperature $T$. While the error bars of the magnetization decrease with the linear system size $L$ for the selfish-energy Glauber dynamics, they do not show a clear system-size dependence near the critical point for constrained Glauber dynamics and Langevin dynamics. The remaining parameter values are the same as in Fig. 3.
E.2 Effects of system size
In this subsection, we investigate the effects of system size as shown in Fig. 7. The average magnetization and specific heat from both Langevin dynamics and the individual-update constrained Glauber dynamics that we propose here exhibit similar behavior across ordered and disordered phases at different system sizes, differing significantly from the results of selfish-energy Glauber dynamics proposed in previous works [30, 29, 50].
For the selfish-energy Glauber dynamics, the error bars of the magnetization decrease as the system size $L$ increases. A similar trend is observed for constrained Glauber dynamics and Langevin dynamics at temperatures away from the critical point. However, near the critical point, no clear correlation between the error bars and the system size is observed.
Figure 8: Average magnetization $\langle m\rangle$ (a) and specific heat $C_{v}$ (b) obtained from Monte-Carlo simulations of Glauber dynamics without auxiliary spins as functions of temperature $T$. The critical temperature is $T_{c}=1.72J$, which differs significantly from the value $T_{c}=0.79J$ obtained through the constrained Glauber dynamics. Parameter values are the same as in Fig. 3.
E.3 Glauber dynamics without auxiliary spins
To show the key role of the auxiliary spins in our Hamiltonian embedding construction, we also consider the Glauber dynamics for the Hamiltonian without auxiliary DOF
$H_{SS}=-\sum_{\langle ij\rangle}\left[J_{ij}(\theta_{i})+J_{ji}(\theta_{j})\right]\cos(\theta_{i}-\theta_{j}).$ (104)
Here, $J_{ij}(\theta_{i})$ is the vision-cone interaction introduced in the Main Text.
The average magnetization and the specific heat are plotted in Fig. 8. The critical temperature is found at $T_{c}=1.72J$, which differs significantly from the value of $T_{c}=0.79J$ obtained from either the constrained Glauber dynamics or the original Langevin dynamics. This result exemplifies the crucial role played by the auxiliary degrees of freedom and the constraint in the Hamiltonian construction proposed in this work. Finally, consistent with the results of Ref. [57], the critical temperature for the Hamiltonian in Eq. (104) is nearly twice that of the value $T_{c}=0.86J$ found for the selfish-energy Glauber dynamics.
Figure 9: Average magnetization $\langle m\rangle$ in the steady state on a square lattice of linear dimension $L=100$, as a function of temperature $T$ for individual-update constrained Glauber (blue dots), selfish-energy Glauber (red dots), and Langevin (green squares) dynamics. The results indicate that in the long-time steady-state, the constrained Glauber dynamics effectively simulates the behavior of the original Langevin dynamics; the agreement is particularly good away from the critical temperature $T_{c}=0.54J$. Close to the critical point, fluctuations prevent reaching a steady state within the $6\times 10^{6}$ ($20\times L^{4}$) iterations used for Langevin (Constrained Glauber), but the critical temperatures agree within the statistical error bars $\Delta T_{c}=0.02J$ (b). By contrast, selfish-energy Glauber dynamics results in significant differences compared to Langevin dynamics, both in the critical region and away from it (a). We use the vision-cone interaction strength in Eq. (105) with $\sigma=1$, along with the semi-width $\Delta=0.5$ for constrained Glauber dynamics, and a time step of $\delta t=0.1$ for the Langevin dynamics.
E.4 Comparison between selfish energy and constrained Glauber dynamics for an alternative vision-cone interaction
To demonstrate that the agreement between the constrained Glauber dynamics and the Langevin dynamics is not a coincidence, we now adopt the vision-cone interaction proposed in Ref. [45] and apply it to the XY spins on the square lattice. Specifically, we consider the interaction strength
$J_{ij}(\theta_{i})=\exp{[\sigma\cos{(\theta_{i}-\psi_{ij})}]},$ (105)
where $\psi_{ij}$ is defined in the main text. The explicit form of the selfish energy [30, 29, 50] for this interaction reads as
$E_{i}=-\sum_{j\in\langle ij\rangle}\exp{[\sigma\cos{(\theta_{i}-\psi_{ij})}]}\cos(\theta_{i}-\theta_{j}).$ ) . (106)
The form of the Hamiltonian embedding remains identical to Eq. (72) with the interaction replaced by Eq. (105). The Langevin equations of motion derived from the Hamiltonian under the constraint $\theta_{i}(t)-\varphi_{i}(t)=\pi$ are given by
$\dot{\theta_{i}}=-\sum_{j\in\langle ij\rangle}\exp{[\sigma\cos{(\theta_{i}-\psi_{ij})}]}\sin(\theta_{i}-\theta_{j})+\sqrt{2T}\eta_{i}(t),$
$\dot{\varphi_{i}}=\sum_{j\in\langle ij\rangle}\exp{[\sigma\cos{(\theta_{i}-\psi_{ij})}]}\sin(\varphi_{i}-\varphi_{j})+\sqrt{2T}\eta_{i}(t).$ ) . (107)
The average steady-state magnetization for the three alternative approaches at system size $L=100$ is shown in Fig. 9, for $\sigma=1$. For the constrained Glauber dynamics, we use semi-width $\Delta=0.5$ and $20\times L^{4}$ iterations; A time step $\delta t=0.1$ is used for Langevin dynamics.
As shown in Fig. 9, the average steady-state magnetization from selfish-energy Glauber dynamics (blue dots) is consistently higher than that obtained from constrained Glauber (red dots) and Langevin (green squares) dynamics. Importantly, the transition temperature from the ordered to the disordered state in the selfish-energy Glauber dynamics differs significantly from that of the other two dynamics for the chosen parameters. Despite the large system size $L=100$, fluctuations in the dynamics of all three approaches remain substantial in the critical region, suggesting that the system has not yet reached a steady state, possibly due to a critical slowing down.
However, we observe that in both the ordered and disordered phases, the constrained Glauber dynamics accurately describes the average magnetization of the Langevin dynamics. In the critical region, there are clearly visible deviations between Langevin dynamics and constrained Glauber dynamics.